### v1

# Triplet scalar flavored leptogenesis with spontaneous CP violation

Abstract
The inclusion of two triplet scalars in the Standard Model (SM) enables to accommodate neutrino mass generation as well as baryogenesis through leptogenesis. One of the essential ingredients of leptogenesis is the violation of charge conjugation and parity (CP) symmetry in lepton number violating decays of the triplet scalars. We work on the promising sector of spontaneous CP violation (SCPV) which is manifested by the involvement of one scalar singlet and two scalar fields, added to the SM. The predictive aspect of the model is accomplished by imposing $A_{4}\times Z_{4}$ symmetry which results in the traditional tribimaximal mixing pattern. With updated data on neutrino oscillation, we study the parameter space of the model. The phase of the complex vacuum expectation value (VEV) of the singlet scalar acts as the common source of CP violation in both low and high energy sectors. Due to the flavor symmetry of the model, required baryon asymmetry cannot be accomplished via unflavored leptogenesis. In the temperature regime, $\left[10^{9},10^{12}\right]$ GeV when flavor effects become important in the study of leptogenesis, it is shown that baryogenesis is achievable. The rich flavor interplay is explored through the study of the density matrix equations. We also study the interplay of hierarchical branching ratios of the decay of the triplet scalars and SCPV phase to accommodate the required CP asymmetry to account for the final baryon asymmetry in the observational range. Considering all possible mass hierarchies among the triplet scalars, the flavor structure of the triplet Yukawa couplings results in different scales of leptogenesis.
1 Introduction
The observations from Big Bang nucleosynthesis (BBN) and cosmic microwave background (CMB) quantify the baryon asymmetry of the Universe (BAU) [46, 35, 32, 23, 5, 6], $\eta_{B}=\frac{(n_{B}-n_{\overline{B}})}{n_{\gamma}}$ as $4.7\times 10^{-10}\leq\eta_{B}\leq 6.5\times 10^{-10}$, where $n_{B}$, $n_{\overline{B}}$, and $n_{\gamma}=\frac{2\zeta(3)}{\pi^{2}}T^{3}$ denote the number densities of baryons, antibaryons and gamma photons, respectively. Baryogenesis through leptogenesis [25] has become a popular theory to explain the matter-antimatter asymmetry of the Universe. Concurrently, thermal leptogenesis could be an important consequence of neutrino mass generation via the seesaw mechanism, since the heavy beyond Standard Model (BSM) particles introduced through the seesaw mechanism could be promising candidates for generating lepton asymmetry. This is how a seesaw framework could explain tiny non-zero neutrino mass and the consequent leptogenesis leading to adequate baryon asymmetry. Apart from the widely studied type-I seesaw mechanism with heavy right-handed neutrinos, type-II and type-III seesaw mechanism with scalar and fermion triplets, respectively, arise naturally from Weinberg’s dimension-5 operator of neutrino mass generation. In this paper, we study baryogenesis through flavored leptogenesis from the decay of triplet Higgs scalar with spontaneous CP violation along with light neutrino mass generated via type-II seesaw mechanism.
Neutrino mass generation through type-II seesaw mechanism [40, 37, 41, 19] and leptogenesis have widely been studied for their unique features coming from the out-of-equilibrium decay of the heavy triplet scalars [30, 39, 28, 18]. Although the addition of just one triplet scalar to the SM can effectively explain the generation of neutrino mass, for producing non-vanishing CP asymmetry for leptogenesis only one triplet scalar is not enough. While a minimal purely triplet leptogenesis can be implemented through two heavy triplet scalars [31, 26, 12], there are more interesting schemes like type-I+II seesaw leptogenesis with one additional right-handed neutrino along with a triplet scalar [27]. In the case of triplet leptogenesis, the flavor interplay comes to be very crucial [36] due to the unique decay channels of the scalar triplets. While flavored leptogenesis has already been studied in the framework of explicit CP violation [20], it is equally appealing to accommodate flavored leptogenesis through SCPV [38, 34, 48, 17, 42]. In this work, we explore the role of the SCPV via the complex VEV of the scalar sector in flavored leptogenesis.
Baryogenesis through leptogenesis from the decay of triplet Higgs has some unique attributes as compared to the other variants of the leptogenesis mechanism. They can be realized from the requirement of CP violation and out-of-equilibrium conditions. One important ingredient of leptogenesis, CP violation, in the case of scalar triplet leptogenesis, can come from two possible sources: (i) Explicit CP violation sourced from the complex Yukawa couplings of the triplet Higgs with leptons and complex scalar coupling to the SM Higgs, (ii) Spontaneous CP violation sourced from the complex VEV of a scalar. The latter one is more economical in terms of parameter counting and attractive from a theoretical point of view [17]. In this case, the CP violation necessary for leptogenesis and potentially observable low-energy leptonic CP violation come from a common source. Also, in models of spontaneously broken CP symmetry, extended by triplet scalars, specific to our case, CP symmetry is imposed at the Lagrangian level, making the Yukawa couplings real. In this case, spontaneous CP violation can be realized through the complex VEVs developed by the triplet scalars of the model. The light neutrino mass matrix being determined by the complex VEV of the triplets and real Yukawa matrices, CP violation can be established at the low scale (such as neutrino oscillation experiments). The high energy CP violation required for the leptogenesis mechanism cannot be accommodated, in this case. However, the model can be extended by singlet scalars to account for both low and high energy CP violation [17]. Some recent studies by adding scalar singlets to the SM that can give rise to spontaneous CP violation and light neutrino mass have been performed in Ref. [16] in the context of dark matter and in Ref. [15], in the context of leptogenesis in vanilla type-I seesaw.
The coupling of the triplet scalars to the leptons and the SM Higgs scalar leads to two different decay channels of the triplet scalar and the lepton number is violated only if both types of decays co-exist. The interplay of the strength of the respective branching ratios, $B^{L}$ and $B^{\phi}$ plays an important role in successful baryogenesis through leptogenesis, where $B^{L}(B^{\phi})$ corresponds to the branching ratios of the triplet scalar to the leptons (SM Higgs scalar). The amount of CP asymmetry, $\epsilon$, and the efficiency of leptogenesis depend on $B^{L}$ and $B^{\phi}$ in a different manner. The efficiency is maximal when $B^{L}\gg B^{\phi}$ or $B^{\phi}\gg B^{L}$, whereas this condition corresponds to suppressed CP violation (as $\epsilon\propto\sqrt{B^{L}B^{\phi}}$). Solving a full set of Boltzmann equations, in a model-independent way, it was shown in Ref. [31] that even for extremely hierarchical branching ratios, it is possible to achieve the required BAU. In this work, we focus on leptogenesis from the decay of triplet scalars with hierarchical branching ratios and CP violation originating from the VEV of a scalar singlet, embedded in the SM along with with two triplet scalars. Also to make the model more predictive we impose $A_{4}$ symmetry at the Lagrangian level. Since in our case, adequate baryon asymmetry is obtained for the triplet scalar mass around $\sim 10^{10}$ GeV, flavor effects become important in this regime. Also, considering hierarchical branching ratios of triplet scalar decays, washout effects play an important role in generating adequate lepton/baryon asymmetry. With the improved computation of the kinetic evolution of flavored lepton asymmetries, we use Density Matrix Equations (DME) for the evaluation of final lepton asymmetries. We also study, with all possible mass hierarchies among the triplet scalars, the flavor structure of the model predicts different scales of leptogenesis.
The paper is organized as follows. The model is introduced in section (2). Section (3) is devoted to pointing out the important features of flavored CP asymmetry and DME for the analysis. Finally, the results are presented in section (4) with the summary. The conclusion is given in section (5). Three appendices (A), (B), (C) are added containing some of the detailed calculations.
Field
$L$
$e_{R},\mu_{R},\tau_{R}$
$\Delta_{1}$
$\Delta_{2}$
$\phi$
$S$
$\Phi$
$\psi$
$SU(2)_{L}$
$2$
$1$
$3$
$3$
$2$
$1$
$1$
$1$
$U(1)_{Y}$
$-1/2$
$-1$
$1$
$1$
$1/2$
00
00
00
$A_{4}$
$3$
$1,1^{\prime\prime},1^{\prime}$
$1$
$1$
$1$
$1$
$3$
$3$
$Z_{4}$
$i$
$-i$
$1$
$-1$
$i$
$-1$
$i$
$1$
Table 1: Field representations under $SU(2)_{L}$, $U(1)_{Y},A_{4}$ and $Z_{4}$ symmetries.
2 Model Description
In this section, we study an extension of the SM, developed in Ref. [17] to account for SCPV. With updated data from neutrino oscillation and cosmology we explore the parameter space of the model. In this model, the SM is extended with one complex scalar singlet $S$ and two Higgs triplet scalars $\Delta_{a}(a=1,2)$. To generate a realistic lepton mixing pattern, two scalar fields, $\Phi$ and $\psi$ are added by imposing $A_{4}$ flavor symmetry and $Z_{4}$ discrete symmetry. The appropriate charge assignments to all the scalar and fermion fields are summarized in table (1). The two triplet Higgs fields can be represented in $SU(2)$ representation as,
$\Delta_{a}=\begin{pmatrix}\Delta^{0}_{a}&-\frac{\Delta^{+}_{a}}{\sqrt{2}}\\ -\frac{\Delta^{+}_{a}}{\sqrt{2}}&\Delta^{++}_{a}\end{pmatrix},$ (2.1)
The product rule to construct singlets from different products of irreducible representations of $A_{4}$ in Altarelli- Feruglio basis [10, 9] is given as
$(a\otimes b)_{1}=a_{1}b_{1}+a_{2}b_{3}+a_{3}b_{2},$
$(a\otimes b)_{1^{\prime}}=a_{3}b_{3}+a_{1}b_{2}+a_{2}b_{1},$
$(a\otimes b)_{1^{\prime\prime}}=a_{2}b_{2}+a_{1}b_{3}+a_{3}b_{1},$ (2.2)
$(a\otimes b)_{3_{s}}=\frac{1}{3}(2a_{1}b_{1}-a_{2}b_{3}-a_{3}b_{2},2a_{3}b_{3}-a_{1}b_{2}-a_{2}b_{1},2a_{2}b_{2}-a_{1}b_{3}-a_{3}b_{1}),$
$(a\otimes b)_{3_{a}}=\frac{1}{2}(a_{2}b_{3}-a_{3}b_{2},a_{1}b_{2}-a_{2}b_{1},a_{1}b_{3}-a_{3}b_{1}),$
$((a\otimes b)_{3_{s}}\otimes c)_{1}$ $=$ $\frac{1}{3}(2a_{1}b_{1}c_{1}+2a_{2}b_{2}c_{2}+2a_{3}b_{3}c_{3}-a_{1}b_{2}c_{3}-a_{1}b_{3}c_{2}$ (2.3)
-- $a_{3}b_{1}c_{2}-a_{3}b_{2}c_{1}-a_{2}b_{1}c_{3}-a_{2}b_{3}c_{1}).$
CP invariance is imposed at the Lagrangian level and hence, all the parameters are perceived to be real. The CP asymmetry is broken at high energy by the complex VEV of the scalar singlet. Also $A_{4}\times Z_{4}$ symmetry is broken at a high-scale with a specific vacuum configuration leading to the tribimaximal (TBM) lepton mixing. Below the cutoff scale, $\Lambda$, the effective Yukawa Lagrangian containing the lowest-order term ($\mathcal{O}(1/\Lambda)$) can be written as,
$\mathcal{L}=\frac{y^{l}_{e}}{\Lambda}(\overline{L}\Phi)_{1}\phi e_{R}+\frac{y^{l}_{\mu}}{\Lambda}(\overline{L}\Phi)_{1^{\prime}}\phi\mu_{R}+\frac{y^{l}_{\tau}}{\Lambda}(\overline{L}\Phi)_{1^{\prime\prime}}\phi\tau_{R}+\frac{y_{2}}{\Lambda}\Delta_{2}(L^{T}L\psi)_{1}+\frac{1}{\Lambda^{\prime}}\Delta_{1}(L^{T}L)_{1}(y_{1}S+y^{\prime}_{1}S^{*})+{\rm h.c.}.$ (2.4)
The heavy scalar fields develop VEVs along the required directions as
$\langle\Phi\rangle=(r,0,0),\quad\langle\psi\rangle=(s,s,s),\quad\langle S\rangle=v_{S}e^{i\alpha}.$ (2.5)
Now the Lagrangian can be written as
$\mathcal{L}$ $=$ $\frac{y^{l}_{e}r}{\Lambda}\overline{L}_{e}\phi e_{R}+\frac{y^{l}_{\mu}r}{\Lambda}\overline{L}_{\mu}\phi\mu_{R}+\frac{y^{l}_{\tau}r}{\Lambda}\overline{L}_{\tau}\phi\tau_{R}+\frac{y_{2}s}{3\Lambda}\Delta_{2}(2L^{T}_{e}L_{e}+2L^{T}_{\mu}L_{\mu}+2L^{T}_{\tau}L_{\tau}-L^{T}_{e}L_{\mu}$ (2.6)
-- $L^{T}_{e}L_{\tau}-L^{T}_{\mu}L_{e}-L^{T}_{\mu}L_{\tau}-L^{T}_{\tau}L_{e}-L^{T}_{\tau}L_{\mu})$
$+$ $\frac{v_{S}}{\Lambda^{\prime}}\Delta_{1}(L^{T}_{e}L_{e}+L^{T}_{\mu}L_{\tau}+L^{T}_{\tau}L_{\mu})(y_{1}e^{i\alpha}+y^{\prime}_{1}e^{-i\alpha})+{\rm h.c.}$
$=$ $Y^{l}_{\alpha\beta}\overline{L}_{\alpha}\phi e_{R\beta}+Y^{\Delta_{1}}_{\alpha\beta}L^{T}_{\alpha}C\Delta_{1}L_{\beta}+Y^{\Delta_{2}}_{\alpha\beta}L^{T}_{\alpha}C\Delta_{2}L_{\beta}+{\rm h.c.},$
where the Yukawa matrices are
$Y^{l}=\begin{pmatrix}y_{e}&0&0\\ 0&y_{\mu}&0\\ 0&0&y_{\tau}\end{pmatrix},\quad Y^{\Delta_{1}}=y_{\Delta_{1}}\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix},\quad Y^{\Delta_{2}}=\frac{y_{\Delta_{2}}}{3}\begin{pmatrix}2&-1&-1\\ -1&2&-1\\ -1&-1&2\end{pmatrix},$ (2.7)
and
$y_{e,\mu,\tau}=\frac{r}{\Lambda}y^{l}_{e,\mu,\tau},\quad y_{\Delta_{1}}=\frac{v_{S}}{\Lambda^{\prime}}(y_{1}e^{i\alpha}+y^{\prime}_{1}e^{-i\alpha}),\quad y_{\Delta_{2}}=\frac{y_{2}}{\Lambda}s.$ (2.8)
It can be seen that the Yukawa coupling matrices $Y^{\Delta_{1}}$ and $Y^{\Delta_{2}}$ exhibit $\mu-\tau$ symmetry and magic symmetry, respectively. We assume the singlet scalar $S$ to be very heavy and to have decoupled at a higher energy scale than electroweak (EW) and mass scales of the Higgs triplets. The complex scalar $S$ plays a vital role in this model since the CP symmetry, otherwise conserved in the Lagrangian level, is spontaneously broken by the complex VEV of the scalar singlet. Hence, it is crucial to analyze the scalar potential for $S$. The detail of the analysis can be found in Appendix (A). The minimization of the potential leads to the following conditions,
$v^{2}_{S}=\frac{-2\lambda^{\prime}_{S}m^{2}_{S}+\lambda^{\prime\prime}_{S}\mu^{2}_{S}}{4\lambda_{S}\lambda^{\prime}_{S}-8\lambda^{\prime}_{S}-\lambda^{\prime\prime}_{S}},\quad\cos(2\alpha)=-\frac{\mu^{2}_{S}+\lambda^{\prime\prime}_{S}v^{2}_{S}}{4\lambda^{\prime}_{S}v^{2}_{S}}.$ (2.9)
To show that the last solution, which also leads in this case to the spontaneous breaking of the CP symmetry, corresponds to the global minimum of the potential, for a particular choice of parameter space such as $m^{2}_{S}<0$, $\lambda^{\prime\prime}_{S}\simeq 0$, $\mu_{S}\simeq 0$ and $\lambda_{S}>2\lambda^{\prime}_{S}>0$ and we obtain from Eq.(A.13), for case-3
$v^{2}_{S}\simeq-\frac{m^{2}_{S}}{2(\lambda_{S}-2\lambda^{\prime}_{S})},\quad\alpha\simeq\pm\frac{\pi}{4},$ (2.10)
which leads to
$V_{0}\simeq-\frac{m^{4}_{S}}{4(\lambda_{S}-2\lambda^{\prime}_{S})}$ (2.11)
which is the absolute minimum of the potential.
The triplet Higgs scalars contribute to the neutrino mass generation through the type-II seesaw mechanism as
$m_{\nu}=m^{(1)}_{\nu}+m^{(2)}_{\nu},\quad m^{(a)}_{\nu}=2u_{a}Y^{\Delta_{a}},$ (2.12)
where $m^{(1)}_{\nu}$ and $m^{(2)}_{\nu}$ are the contributions coming from the triplet scalars $\Delta_{1}$ and $\Delta_{2}$, respectively. Here, $u_{a}$ is the VEV of the neutral components of the triplet scalars $\Delta_{a}$, and it is given by
$u_{a}=\mu^{*}_{a}v^{2}/M_{a},\quad\mu_{1}=(\lambda_{1}e^{i\alpha}+\lambda^{\prime}_{1}e^{-i\alpha})v_{S}/M_{1}$ (2.13)
where $v=\langle\phi_{0}\rangle=174$ GeV. To diagonalize the neutrino mass matrix, we define a unitary matrix $U$ as
$m_{\nu}=U^{*}d_{\nu}U^{\dagger},\quad d_{\nu}=\rm{diag}(|z_{1}e^{i\beta}+z_{2}|,z_{1},|z_{1}e^{i\beta}-z_{2}|)\equiv diag(m_{1},m_{2},m_{3}),$ (2.14)
where
$z_{a}=2|u_{a}y_{\Delta_{a}}|,\quad\beta={\rm arg}(u_{1}y_{\Delta_{1}})={\rm arctan}\left(\frac{\lambda^{\prime}_{1}-\lambda_{1}}{\lambda^{\prime}_{1}+\lambda_{1}}{\rm tan}\alpha\right)+{\rm arctan}\left(\frac{y^{\prime}_{1}-y_{1}}{y^{\prime}_{1}+y_{1}}{\rm tan}\alpha\right).$ (2.15)
In this model, where the neutrino mass matrix can be decomposed with matrices with $\mu-\tau$ symmetry and magic symmetry, the mixing matrix takes the form
$U=e^{-i\sigma_{1}/2}U_{\rm{TBM}}K,$ (2.16)
where $U_{\rm{TBM}}$ is the tribimaximal mixing matrix and is given by,
$U_{\rm{TBM}}=\begin{pmatrix}\frac{2}{\sqrt{6}}&\frac{1}{\sqrt{3}}&0\\ -\frac{1}{\sqrt{6}}&\frac{1}{\sqrt{3}}&-\frac{1}{\sqrt{2}}\\ -\frac{1}{\sqrt{6}}&\frac{1}{\sqrt{3}}&\frac{1}{\sqrt{2}}\end{pmatrix},$ (2.17)
and the Majorana phases are the entries in the diagonal matrix $K$, given by,
$K=\rm{diag}(1,e^{i\gamma_{1}},e^{i\gamma_{2}}),$ (2.18)
with
$\gamma_{1}=(\sigma_{1}-\beta)/2,\quad\gamma_{2}=(\sigma_{1}-\sigma_{2})/2,\quad\sigma_{1,2}=\rm{arg}(z_{2}\pm z_{1}e^{i\beta}).$ (2.19)
TBM mixing matrix has gathered considerable attention because of its proximity with the observation of neutrino masses and mixing. However the absence of Dirac-type CP violating phase asks for possible modifications of the model, which we discuss in the next subsection. Considering the two cases of neutrino mass hierarchy, we can have two sets of constraints on $z_{1}$ and $z_{2}$. From Eq.(2.14), for normal hierarchy (NH), i.e. $m_{3}>m_{2}>m_{1}$, we obtain
$z_{2}+2z_{1}\cos\beta<0,\quad z_{2}-2z_{1}\cos\beta>0,$ (2.20)
and, for inverted hierarchy (IH) we have
$z_{2}+2z_{1}\cos\beta<0,\quad z_{2}-2z_{1}\cos\beta<0.$ (2.21)
Now the constraints in Eq.(2.21) cannot be satisfied simultaneously, which indicates that the model under consideration does not qualify for the IH of neutrino mass. In terms of the phase $\beta$ and two mass squared differences, the parameters $z_{1}$ and $z_{2}$ can be written as,
$z_{1}=-\frac{1}{2\cos\beta}\frac{\Delta m^{2}_{31}}{\sqrt{2(\Delta m^{2}_{31}-2\Delta m^{2}_{21})}}\simeq-\frac{1}{2\cos\beta}\sqrt{\frac{\Delta m^{2}_{31}}{2}},$ (2.22)
and
$z_{2}=\sqrt{\frac{\Delta m^{2}_{31}-2\Delta m^{2}_{21}}{2}}\simeq\sqrt{\frac{\Delta m^{2}_{31}}{2}}.$ (2.23)
Hence, the neutrino mass eigenvalues can be expressed as
$m_{1}=\sqrt{z^{2}_{1}-\Delta m^{2}_{21}},\quad m_{2}=z_{1},\quad m_{3}=\sqrt{z^{2}_{1}+\Delta m^{2}_{31}-\Delta m^{2}_{21}}.$ (2.24)
2.1 Non-zero $\theta_{13}$ from perturbative corrections in the flavon VEV
The measurement of non-zero $\theta_{13}$ [1, 4, 11] requires to modify the TBM scenario. It can be achieved by exploiting the small perturbations around the TBM vacuum alignment conditions. More specifically, by considering a small perturbation around the flavon VEV of the form $\langle\Phi\rangle=r(1,\varepsilon_{1},\varepsilon_{2})$ with $|\varepsilon_{1,2}|\ll 1$ we obtain the transformed charged lepton Yukawa matrix,
$Y^{l}=\begin{pmatrix}y_{e}&y_{\mu}\varepsilon_{1}&y_{\tau}\varepsilon_{2}\\ y_{e}\varepsilon_{2}&y_{\mu}&y_{\tau}\varepsilon_{1}\\ y_{e}\varepsilon_{1}&y_{\mu}\varepsilon_{2}&y_{\tau}\end{pmatrix},$ (2.25)
while the lepton mixing coming from the neutrino sector remains invariably of TBM type, since $\langle\psi\rangle$ is not perturbed. The emergence of a non-diagonal charged lepton Yukawa matrix can be diagonalized by biunitary transformation $Y^{l}=U_{l}~{}{\rm Diagonal}(y_{e},y_{\mu},y_{\tau})~{}I$, where $I$ is the identity matrix and $U_{l}$ is given by
$U_{l}=\begin{pmatrix}1&\varepsilon_{1}&\varepsilon_{2}\\ \varepsilon_{2}&1&\varepsilon_{1}\\ \varepsilon_{1}&\varepsilon_{2}&1\end{pmatrix}.$ (2.26)
Further, the lepton mixing matrix gets modified as
$U=e^{-i\sigma_{1}/2}U^{\dagger}_{l}U_{TBM}K.$ (2.27)
The $U$ matrix obtained in Eq. (2.27) can be identified as Pontecorvo-Maki-Nakagawa-Sakata (PMNS) mixing matrix, $U_{{\rm PMNS}}$.
2.2 Parameter space study
To check the compatibility of the model with the latest experimental results, it is important to study the parameter space of the model constrained by the neutrino oscillation data. The latest global-fit of the observational data are summarized in table (2) [24]. Also, the limits for effective mass, $|m_{ee}|$ appearing in the half-life of neutrinoless double beta ($0\nu\beta\beta$) decay are constrained by GERDA as $<(79-180)$ meV [7] and KamLAND-Zen as $<(36-156)$ meV [2]. The future sensitivity of $|m_{ee}|$ from the future ton-scale experiments like CUPID [14, 13], LEGEND [3], and nEXO [8, 33] are going to probe $|m_{ee}|\leq 0.01$ eV.
NuFIT 5.0 [24]
Neutrino Oscillation parameters
[…]
$\sin^{2}\theta_{12}$
$0.303^{+0.012}_{-0.012}$
$0.270\rightarrow 0.341$
$\sin^{2}\theta_{23}$
$0.451^{+0.019}_{-0.016}$
$0.408\rightarrow 0.603$
$\sin^{2}\theta_{13}$
$0.02225^{+0.00056}_{-0.00059}$
$0.02052\rightarrow 0.02398$
$\delta/^{\circ}$
$232^{+36}_{-26}$
$144\rightarrow 350$
$\frac{\Delta m^{2}_{21}}{10^{-5}{\rm eV}^{2}}$
$7.41^{+0.21}_{-0.20}$
$6.82\rightarrow 8.03$
$\frac{\Delta m^{2}_{3l}}{10^{-3}{\rm eV}^{2}}$
$+2.507^{+0.026}_{-0.027}$
$+2.427\rightarrow+2.590$
Table 2: Allowed ranges of neutrino oscillation parameters for normal mass ordering.
The three mixing angles, $\theta_{12}$, $\theta_{23}$ and $\theta_{13}$ can be extracted from the standard parameterization of the $U_{\rm PMNS}$ matrix as
$\sin^{2}\theta_{12}=\frac{|U_{12}|^{2}}{1-|U_{13}|^{2}},\quad\sin^{2}\theta_{23}=\frac{|U_{23}|^{2}}{1-|U_{13}|^{2}},\quad\sin^{2}\theta_{13}=|U_{13}|^{2}.$ (2.28)
From equations (2.27) and (2.28), three mixing angles, $\theta_{12}$, $\theta_{23}$ and $\theta_{13}$ are related to the small perturbations $\varepsilon_{1}$ and $\varepsilon_{2}$ through the $U$ matrix, given in Eq. (2.27). The explicit relations relating the three mixing angles and the small perturbations are shown below:
$\sin^{2}\theta_{12}\simeq\frac{1}{3}[1+2(\varepsilon_{1}+\varepsilon_{2})],\quad\sin^{2}\theta_{23}\simeq\frac{1}{2}[1-2\varepsilon_{2}],\quad\sin^{2}\theta_{13}\simeq\frac{1}{2}(\varepsilon_{1}-\varepsilon_{2})^{2},$ (2.29)
which further satisfies the relation,
$\sin^{2}\theta_{13}\simeq\frac{(4\sin^{2}\theta_{23}-3\cos^{2}\theta_{12})^{2}}{8}.$ (2.30)
Figure 1: Left: The plot shows the parameter space for the perturbations $\varepsilon_{1}$ and $\varepsilon_{2}$ constrained by the $3\sigma$ data of three mixing angles. Right: The plot shows the dependence of $|m_{ee}|$ (eV) on the phase $\beta$ (rad) using the bounds of the $3\sigma$ data of three mixing angles and the two mass squared differences. The experimental limits on $|m_{ee}|$ are given by GERDA (cyan) and KamLAND-Zen (magenta). The orange line corresponds to the future sensitivity of $|m_{ee}|$ from the future ton-scale experiments like CUPID, LEGEND, and nEXO.
By imposing the experimentally obtained constraints on the mixing angle, as given in table (2), the allowed parameter space for the perturbations $[\varepsilon_{1},\varepsilon_{2}]$ are obtained as shown in Fig. (0(a)). We use the same set of mixing angle constraints along with the bounds on two mass squared differences to assess the effective neutrino mass, which is given by
$|m_{ee}|=\frac{1}{3}\left|2m_{1}(1-\varepsilon_{1}-\varepsilon_{2})+m_{2}(1+2\varepsilon_{1}+2\varepsilon_{2})e^{2i\gamma_{1}}\right|,$ (2.31)
where the phase $\gamma_{1}$ is given in Eq. (2.19). Through the phase, $\gamma_{1}$, it is possible to draw a link between the effective mass and its dependence on the phase $\beta$, which appears in the expressions of neutrino masses as well as in the high-scale CP violation, as it will be elaborated later on. In Fig. (0(b)) we show the dependence of $|m_{ee}|$ on the phase $\beta$. The plot shows that the allowed effective mass in turn has imposed a bound on the phase $\beta\leq 270^{\circ}$. Apart from the bound on the effective neutrino mass $|m_{ee}|\lesssim 0.15$ eV, the bound on the total neutrino mass $\Sigma=m_{1}+m_{2}+m_{3}<0.12$ eV [6] puts a more stringent restriction on the parameter space of phase $\beta$ as $\sim[120^{\circ},240^{\circ}]$. We have used this restricted parameter space of the phase $\beta$ to calculate the flavored CP asymmetries in section (4).
3 Baryogenesis through leptogenesis
The Sakharov conditions [44] provide necessary ingredients for dynamical generation of BAU viz; baryon number violation, C and CP violation and out-of-equilibrium dynamics in the processes involving baryons. In the framework of baryogenesis through leptogenesis, the necessary CP violation required for successful leptogenesis is provided by the decay of heavy particles added to the SM. In the type-II seesaw scenario, the decays of the triplet scalars $\Delta_{1}$ and $\Delta_{2}$ serve the purpose of generating adequate CP asymmetry.
The flavor-independent CP asymmetry, as shown in Eq. (3.9), crucially depends on the factor ${\rm Tr}(m^{(a)}_{\nu}m^{(b)\dagger}_{\nu})$ that comes from the neutrino mass matrices given in Eq. (2.12). In the framework under consideration, this factor vanishes due to the specific flavor structure of the Yukawa matrices $Y^{\Delta_{a}}$, causing the flavor-independent CP asymmetry to diminish. Hence, the flavor-independent analysis of leptogenesis is not feasible in our chosen framework. Moreover, since scalar triplets generally couple to lepton pairs from different flavor generations, it is more appropriate to consider the flavor effects while studying triplet scalar leptogenesis.
3.1 CP violation
A triplet scalar can decay into two leptons or two SM Higgs doublets at tree level. Lepton asymmetry can be produced only if the two decay channels co-exist. The non-vanishing lepton asymmetry can be produced for each triplet component from the decay of triplet scalar $\Delta_{a}$ into two leptons and the CP asymmetry can be estimated through,
$\epsilon_{a}=\Delta L\sum_{\alpha\beta}\frac{\Gamma(\Delta^{*}_{a}\rightarrow L_{\alpha}+L_{\beta})-\Gamma(\Delta_{a}\rightarrow\overline{L}_{\alpha}+\overline{L}_{\beta})}{\Gamma_{\Delta_{a}}+\Gamma_{\Delta^{*}_{a}}},$ (3.1)
where $\Delta L=2$ since the lepton number is violated by two units. The tree level branching ratios $B^{L}_{a}$ and $B^{\phi}_{a}$,
$B^{L}_{a}\Gamma_{\Delta_{a}}\equiv\sum_{\alpha,\beta}\Gamma(\Delta^{*}_{a}\rightarrow L_{\alpha}+L_{\beta})=\frac{M_{a}}{8\pi}{\rm Tr}(Y^{\Delta_{a}\dagger}Y^{\Delta_{a}}),$ (3.2)
$B^{\phi}_{a}\Gamma_{\Delta_{a}}\equiv\Gamma(\Delta^{*}_{a}\rightarrow\phi+\phi)=\frac{M_{a}}{8\pi}|\mu_{a}|^{2},$ (3.3)
satisfy the relation $B^{L}_{a}+B^{\phi}_{a}=1$, provided
$\Gamma_{\Delta_{a}}=\frac{M_{a}}{8\pi}\left[{\rm Tr}(Y^{\Delta_{a}\dagger}Y^{\Delta_{a}})+|\mu_{a}|^{2}\right],$ (3.4)
which is the total triplet decay width. In the case of a lepton flavor violating interaction, the interference of tree level (Fig. 2) and one-loop (Fig. 3) diagrams give rise to a non-vanishing CP asymmetry, given by
$\epsilon^{\alpha\beta}_{a}\simeq-\frac{g(x_{b})}{2\pi}\frac{c_{\alpha\beta}{\rm Im}[\mu^{*}_{a}\mu_{b}Y^{\Delta_{a}}_{\alpha\beta}Y^{\Delta_{b}*}_{\alpha\beta}]}{{\rm Tr}(Y^{\Delta_{a}\dagger}Y^{\Delta_{a}})+|\mu_{a}|^{2}},$ (3.5)
where $x_{b}=\frac{M^{2}_{b}}{M^{2}_{a}}$, and the on-loop self energy function,
$g(x_{b})=\frac{\sqrt{x_{b}}(1-x_{b})}{(x_{b}-1)^{2}+(\Gamma_{\Delta_{b}}/M_{a})^{2}},$ (3.6)
and
$c_{\alpha\beta}$ $=$ $2-\delta_{\alpha\beta}\quad\rm{for}\quad\Delta^{0}_{a},\Delta^{++}_{a},$
$=$ $1\quad\rm{for}\quad\Delta^{+}_{a}.$
In the present scenario, we can rewrite Eq. (3.5) as
$\epsilon^{\alpha\beta}_{a}\simeq-\frac{g(x_{b})}{4\pi}\frac{M_{b}(B^{L}_{a}B^{\phi}_{a})^{\frac{1}{2}}}{v^{2}}\frac{c_{\alpha\beta}{\rm Im}[m^{(a)}_{\nu,\alpha\beta}m^{*}_{\nu,\alpha\beta}]}{[{\rm Tr}(m^{(a)\dagger}_{\nu}m^{(a)}_{\nu})]^{1/2}}.$ (3.7)
If we consider the triplet scalar masses to be hierarchical, satisfying $M_{a}\ll M_{b}$ then
$\epsilon^{\alpha\beta}_{a}\simeq\frac{M_{a}(B^{L}_{a}B^{\phi}_{a})^{\frac{1}{2}}}{v^{2}}\frac{c_{\alpha\beta}{\rm Im}[m^{(a)}_{\nu,\alpha\beta}m^{*}_{\nu,\alpha\beta}]}{[{\rm Tr}(m^{(a)\dagger}_{\nu}m^{(a)}_{\nu})]^{1/2}}.$ (3.8)
We obtain the flavor independent or unflavored CP asymmetry from Eq. (3.8), by summing over all lepton flavors,
$\epsilon_{a}\simeq\frac{M_{a}(B^{L}_{a}B^{\phi}_{a})^{\frac{1}{2}}}{v^{2}}\frac{{\rm Im}[{\rm Tr}(m^{(a)}_{\nu}m^{\dagger}_{\nu})]}{[{\rm Tr}(m^{(a)\dagger}_{\nu}m^{(a)}_{\nu})]^{1/2}}.$ (3.9)
This CP asymmetry has an upper bound as can be seen from the expression,
$|\epsilon_{a}|\leq\frac{M_{a}(B^{L}_{a}B^{\phi}_{a})^{\frac{1}{2}}}{v^{2}}[{\rm Tr}(m^{(a)}_{\nu}m^{\dagger}_{\nu})]^{1/2}=\frac{M_{a}(B^{L}_{a}B^{\phi}_{a})^{\frac{1}{2}}}{v^{2}}\left(\sum_{k}m^{2}_{k}\right)^{1/2}.$ (3.10)
Further, for hierarchical light neutrinos, one finds
$|\epsilon_{a}|\lesssim 10^{-6}(B^{L}_{a}B^{\phi}_{a})^{\frac{1}{2}}\left(\frac{M_{a}}{10^{10}GeV}\right)\left(\frac{\sqrt{\Delta m^{2}_{31}}}{0.05eV}\right)$ (3.11)
which reaches its absolute maxima at $B^{L}_{a}=B^{\phi}_{a}=\frac{1}{2}$.
Figure 2: Tree-level decay diagram: Triplet scalar decays into lepton and SM Higgs, respectively.
Figure 3: One-loop decay diagram: Triplet scalar decays into leptons through Higgs loop and lepton loop, respectively.
3.2 Triplet leptogenesis and flavor covariance
Flavor effects can emerge from the misalignment of flavor and mass eigenbasis, which can persist in the off-diagonal entries of charged lepton Yukawa matrices. The phenomenon of particle mixing with primordial plasma as a background medium also contributes to flavor effects in the early universe. Such effects can be addressed through quantum statistical mechanics, precisely considering the non-vanishing off-diagonal entries of particle number densities of particle species tagged with flavor quantum numbers. In the case of thermal leptogenesis, the charged lepton flavor coherence affects the washout process, thereby affecting the generation of baryon asymmetry. Hence it is necessary to develop theoretical frameworks to assess the flavor effects. We will use a formalism, namely closed time path (CTP) formalism, in which particle densities are replaced by Green’s functions to describe non-equilibrium phenomena like thermal leptogenesis [36, 22, 47, 43, 45]. The flavor covariance could be conveniently studied in a density matrix formalism [36] through flavor-covariant Boltzmann equations. The temperature range $[10^{9},10^{12}]$ GeV appears more interesting for the study of such flavor covariance since this is an intermediate temperature range $\lesssim 10^{12}$ GeV, when the flavor coherence between $\tau$-leptons and the other two leptons $(e,\mu)$ is broken, yet before the complete disintegration of the flavor coherence that happens further below $10^{9}$ GeV.
Hence, we will study the leptogenesis through a set of flavor-covariant Boltzmann equations and perceive the asymmetries in terms of a density matrix $(\Delta_{l})_{\alpha\beta}$. In general, it is a $3\times 3$ matrix in lepton flavor space. The lepton doublet asymmetries are stored as diagonal entries, whereas, the off-diagonal entries symbolise the quantum correlation between the different flavors. In the most general way, we can write such a flavor-covariant equation to express the evolution of flavored lepton asymmetries, as
$sHz\frac{d(\Delta_{l})_{\alpha\beta}}{dz}=\rm{Source}+\rm{Washout}.$ (3.12)
The source term on the right-hand side is proportional to the flavor-covariant CP asymmetry matrix $\epsilon^{\alpha\beta}$. Different lepton number violating processes like inverse triplet and anti-triplet decays washout the lepton asymmetry generated through decay processes. Similarly, $\Delta L=2$ scatterings via s - and t - channel triplet exchange, 2 - lepton - 2 - lepton scatterings, as well as lepton flavor violating processes also contribute to affect the total lepton asymmetry. So, they come under the class of washout processes. To avoid complexity, we will only stick to the washout associated with inverse decays. Apart from the inverse decay and scattering processes, there are different SM reactions, such as strong and EW sphaleron process and Yukawa couplings that indirectly affect the final asymmetry. They are considered spectator processes. For a specific temperature range, these spectator processes in thermal equilibrium can impose some relations among chemical potentials and thereby among the asymmetries. As a consequence, different asymmetries appearing in the Boltzmann equation can be expressed in terms of asymmetries conserved by all active SM processes.
We write the evolution of the asymmetries of the particle species under consideration in a set of flavor-covariant Boltzmann equations, with respect to a factor $z=\frac{M_{a}}{T}$, $T$ is the temperature of the Universe, $H(z)$ is the Hubble rate of expansion of the Universe, $H(z)=\frac{H_{0}(M_{a})}{z^{2}}$, $H_{0}(T)=\sqrt{\frac{4\pi^{3}}{45}g_{*}}\frac{T^{2}}{M_{P}}$, $M_{P}=1.22\times 10^{19}$ GeV is the Planck mass, $s=\left(\frac{2\pi^{2}}{45}\right)g_{*}T^{3}$ is the total entropy density and $g_{*}=106.75$. The asymmetries are defined as $\Delta_{x}=\frac{n_{x}-n_{\overline{x}}}{s}$ where $n_{x}(n_{\bar{x}})$ is the number density of the species, $x(\bar{x})$, and $\Sigma_{\Delta}=\frac{n_{\Delta}+n_{\overline{\Delta}}}{s}$ is the total triplet number density. Also, $Y_{x}=\frac{n_{x}}{s}$ is the comoving number density,
$Y^{\rm eq}_{\Delta}=\frac{45g_{\Delta}}{4\pi^{4}g_{*}}z^{2}K_{2}(z),\quad Y^{\rm eq}_{L}=\frac{3}{4}\frac{45\varsigma(3)}{2\pi^{4}g_{*}}g_{L},\quad Y^{\rm eq}_{\phi}=\frac{45\varsigma(3)}{2\pi^{4}g_{*}}g_{\phi},$
are the equilibrium values, with $g_{\Delta}=1$ for each triplet component, $g_{L}=2$ and $g_{\phi}=2$, $\varsigma(3)\simeq 1.202$, $K_{2}(z)$ is the modified Bessel function. Different reaction densities are elaborated in Appendix (B).
We are interested to study the evolution of the asymmetries $\Delta_{\phi}$, $\Delta_{\Delta}$ and $(\Delta_{l})_{\alpha\beta}$ through a set of flavor-covariant Boltzmann equations. Within the temperature range $T\subset[10^{9},10^{12}]$ GeV, the $\tau$-Yukawa couplings enter equilibrium and consequently, the $(e,\tau)$, $(\mu,\tau)$, $(\tau,e)$, $(\mu,e)$ entries of the $3\times 3$ density matrix $(\Delta_{l})_{\alpha\beta}$ come to be zero. So it is convenient to assign $2\times 2$ matrix $(\Delta^{0}_{l})_{\alpha\beta}$ $(\alpha,\beta=e,\mu)$ and $\Delta_{l_{\tau}}$ instead of the $3\times 3$ density matrix $(\Delta_{l})_{\alpha\beta}$ with $(\alpha,\beta=e,\mu,\tau)$ to express the asymmetries and quantum correlations stored in $e-\mu$ flavored lepton subspace and the asymmetries stored in $\tau$-lepton, respectively. Among all the asymmetries, $\Delta_{\phi}$, $\Delta_{\Delta}$, $\Delta_{l_{\tau}}$, $(\Delta^{0}_{l})_{\alpha\beta}$, only $\Delta_{\Delta}$ is conserved by all the SM interactions. Hence, it is necessary to define conserved asymmetries $\Delta_{\tau}$ and $\Delta^{0}_{\alpha\beta}$ (instead of $\Delta_{l_{\tau}}$ and $(\Delta^{0}_{l})_{\alpha\beta}$) along with $\Delta_{\Delta}$ to be used in the Boltzmann equations. The Higgs and flavored lepton asymmetries can be expressed as functions of these conserved asymmetries ($\Delta_{\Delta}$, $\Delta_{\tau}$, $\Delta^{0}_{\alpha\beta}$) using suitable chemical potential relations (discussed in equations (B.4 - B.10)). For the temperature range $T\subset[10^{9},10^{12}]$ GeV, we obtain the relations as [36]
$(\Delta^{0}_{l})_{\alpha\beta}=\left(\frac{86}{589}{\rm Tr}(\Delta^{0}_{\alpha\beta})+\frac{60}{589}\Delta_{\tau}+\frac{8}{589}\Delta_{\Delta}\right)\delta_{\alpha\beta}-\Delta^{0}_{\alpha\beta},$ (3.13)
$\Delta_{l_{\tau}}=\frac{30}{589}{\rm Tr}(\Delta^{0}_{\alpha\beta})-\frac{390}{589}\Delta_{\tau}-\frac{52}{589}\Delta_{\Delta},$ (3.14)
and
$\Delta_{\phi}=-\frac{164}{589}{\rm Tr}\left(\Delta^{0}_{\alpha\beta}\right)-\frac{224}{589}\Delta_{\tau}-\frac{344}{589}\Delta_{\Delta},$ (3.15)
where $\alpha$, $\beta$ describe any two orthogonal in the $(l_{e},l_{\mu})$ flavor subspace now. Therefore, within the temperature range $T\subset[10^{9},10^{12}]$ GeV, the flavor-covariant Boltzmann equations are given by
$sHz\frac{d\Sigma_{\Delta}}{dz}=-\left(\frac{\Sigma_{\Delta}}{\Sigma^{eq}_{\Delta}}-1\right)\gamma_{D}-2\left(\left(\frac{\Sigma_{\Delta}}{\Sigma^{eq}_{\Delta}}\right)^{2}-1\right)\gamma_{A},$ (3.16)
$sHz\frac{d\Delta^{0}_{\alpha\beta}}{dz}=-\left(\frac{\Sigma_{\Delta}}{\Sigma^{eq}_{\Delta}}-1\right)\gamma_{D}\epsilon^{\alpha\beta}+\tilde{\mathcal{W}}^{D}_{\alpha\beta},$ (3.17)
$sHz\frac{d\Delta_{\tau}}{dz}=-\left(\frac{\Sigma_{\Delta}}{\Sigma^{eq}_{\Delta}}-1\right)\gamma_{D}\epsilon^{\tau\tau}+\tilde{\mathcal{W}}^{D}_{\tau},$ (3.18)
$sHz\frac{d\Delta_{\Delta}}{dz}=-\frac{1}{2}\tilde{\mathcal{W}}^{D}_{\Delta},$ (3.19)
where, $\tilde{\mathcal{W}}^{D}$ represent different washout terms associated with inverse decay processes. In order to avoid complexity of presentation, we represent $Y^{\Delta_{a}}\equiv Y$ and
$\tilde{\mathcal{W}}^{D}_{\alpha\beta}=\frac{2B^{L}}{\lambda^{2}_{l}}\left[(YY^{\dagger})_{\alpha\beta}\frac{\Delta_{\Delta}}{\Sigma_{\Delta}}+\frac{1}{4Y^{eq}_{L}}\left(2Y(\Delta^{0}_{l})^{T}Y^{\dagger}+YY^{\dagger}\Delta^{0}_{l}+\Delta^{0}_{l}YY^{\dagger}\right)_{\alpha\beta}+\frac{1}{2Y^{eq}_{L}}Y_{\alpha\tau}Y^{*}_{\beta\tau}\Delta_{l_{\tau}}\right]\gamma_{D},$ (3.20)
$\tilde{\mathcal{W}}^{D}_{\Delta}=Tr(\tilde{\mathcal{W}}^{D})+\tilde{\mathcal{W}}^{D}_{\tau}-W^{D}_{\phi},$ (3.21)
$\tilde{\mathcal{W}}^{D}_{\tau}=\frac{2B^{L}}{\lambda^{2}_{l}}\left[\left(YY^{\dagger}\right)_{\tau\tau}\frac{\Delta_{\Delta}}{\Sigma^{eq}_{\Delta}}+\frac{1}{2Y^{eq}_{L}}\left(\left(Y(\Delta^{0}_{l})^{T}Y^{\dagger}\right)_{\tau\tau}+\left(\left(YY^{\dagger}\right)_{\tau\tau}+|Y_{\tau\tau}|^{2}\right)\Delta_{l_{\tau}}\right)\right]\gamma_{D},$ (3.22)
and
$W^{D}_{\phi}=2B^{\phi}\left(\frac{\Delta_{\phi}}{Y^{eq}_{\phi}}-\frac{\Delta_{\Delta}}{\Sigma^{eq}_{\Delta}}\right)\gamma_{D},$ (3.23)
where $(\Delta^{0}_{l})_{\alpha\beta}$, $\Delta_{l_{\tau}}$ and $\Delta_{\phi}$ are given by equations (3.13 - 3.15).
Solving the flavor-covariant Boltzmann equations numerically, we obtain flavored lepton asymmetries and we estimate the final baryon asymmetry as [36]
$\eta_{B}=7.04\times\frac{12}{37}\times\sum_{\alpha}\Delta_{\alpha},$ (3.24)
where $\Delta_{\alpha}=\Delta_{B/3-L_{\alpha}}$.
4 Baryogenesis through flavored leptogenesis: Results
We obtain the final baryon asymmetry given in Eq.(3.24) via lepton asymmetries $\Delta_{\alpha}$ by solving the full set of DMEs given in equations (3.16 - 3.19), for two cases. The cases depend upon the mass hierarchy among the two triplet scalars; (I) $M_{1}\ll M_{2}$ (II) $M_{2}\ll M_{1}$. In thermal leptogenesis, the Yukawa couplings of the corresponding triplet scalars in both cases play an important role in the generation of adequate CP asymmetries and washout effects. We further study, in each case, the role of hierarchical branching ratios $(B^{L}_{a}\sim 0.0005,0.005,0.05)$ in generating adequate baryon asymmetry as compared to comparable branching ratios $(B^{L}_{a}\sim 0.5)$. The CP-violating phase $\alpha/\beta$ plays an important role in compensating the suppression in the CP asymmetry, which can manifest due to hierarchy in the branching ratios. We also take different sets of triplet $\Delta_{a}$ masses $M_{a}\sim 10^{10}$ GeV, flavored CP asymmetries $\epsilon^{\alpha\beta}_{a}$ and calculate the baryon asymmetry for different hierarchy of branching ratios.
4.1 Formulation of flavored CP asymmetries
The expression of flavored CP asymmetry from triplet $\Delta_{1}$ decay
$\epsilon^{\alpha\beta}_{1}=c_{\alpha\beta}\frac{M^{2}_{1}|u_{1}|^{2}}{2\pi\left(3z^{2}_{1}v^{4}+4M^{4}_{1}|u_{1}|^{4}\right)}\times\rm{Im}\left[m^{(1)}_{\nu,\alpha\beta}m^{*}_{\nu,\alpha\beta}\right],$ (4.1)
and the same originating from triplet $\Delta_{2}$ decay
$\epsilon^{\alpha\beta}_{2}=c_{\alpha\beta}\frac{M^{2}_{2}|u_{2}|^{2}}{4\pi\left(z^{2}_{2}v^{4}(1-2\varepsilon_{1}-2\varepsilon_{2})+2M^{4}_{2}|u_{2}|^{4}\right)}\times\rm{Im}\left[m^{(2)}_{\nu,\alpha\beta}m^{*}_{\nu,\alpha\beta}\right],$ (4.2)
get generated from Eq. (3.8) using the expressions of branching ratios (Eq. (3.3)) and total triplet decay width (Eq. (3.4)). The total neutrino mass matrix expression (Eq. (2.14)) is obtained by taking the modified diagonalising matrix $U$ given in Eq. (2.27). The $\gamma_{1,2}$, $\beta$ phase relations are taken from Eq. (2.19). Using the neutrino mass matrix relation
$m_{\nu}=U^{*}d_{\nu}U^{\dagger}=U^{*}d^{(1)}_{\nu}U^{\dagger}+U^{*}d^{(2)}_{\nu}U^{\dagger},$ (4.3)
we can obtain the neutrino mass matrix contribution from triplet $\Delta_{1}$, as
$m^{(1)}_{\nu}=U^{*}d^{(1)}_{\nu}U^{\dagger}=z_{1}e^{i\beta}\begin{pmatrix}1&2\varepsilon_{1}&2\varepsilon_{2}\\ 2\varepsilon_{1}&2\varepsilon_{2}&1\\ 2\varepsilon_{2}&1&2\varepsilon_{1}\end{pmatrix}$ (4.4)
where
$d^{(1)}_{\nu}=\rm diag\left(z_{1}e^{i(\beta-\sigma_{1})},z_{1},-z_{1}e^{i(\beta-\sigma_{2})}\right).$ (4.5)
and we can similarly obtain the neutrino mass matrix contribution from triplet $\Delta_{2}$, as
$m^{(2)}_{\nu}=U^{*}d^{(2)}_{\nu}U^{\dagger}=\frac{1}{3}z_{2}(1-\varepsilon_{1}-\varepsilon_{2})\begin{pmatrix}2&-1&-1\\ -1&2&-1\\ -1&-1&2\end{pmatrix}$ (4.6)
$d^{(2)}_{\nu}=\rm diag\left(z_{2},0,z_{2}\right).$ (4.7)
Within the energy range $[10^{9},10^{12}]$ GeV, the flavored CP asymmetries can be expressed as $\epsilon^{\alpha\beta}_{a}(\alpha,\beta=e,\mu)$ and $\epsilon^{\tau\tau}_{a}$, where $a=1,2$ is associated with triplet $\Delta_{1,2}$. Using equations (4.3), (4.4) and (4.6) we calculate, the flavored CP asymmetries. The specific expressions are given in equations (C.8 - C.11) and (C.12 - C.15) in Appendix (C). The dependence of $\beta$ on the CP asymmetries are plotted in the figures (8) and (9) for different values $B^{L}_{1(2)}=0.5,0.05,0.005,0.0005$. The triplet VEV $|u_{1}|$ ($|u_{2}|$) are taken in the range $[0.1-100]$ eV. To estimate the CP asymmetry from the model, we have chosen the small perturbation parameters from the allowed parameter space shown in Fig.(0(a)), as $\varepsilon_{1}\sim-0.125$ and $\varepsilon_{2}\sim 0.075$. We have varied the phase $\beta$ within the range $[120^{\circ},240^{\circ}]$ so that it also remains consistent with the latest cosmological bound on the sum of the neutrino masses and the effective neutrino mass $|m_{ee}|\lesssim 0.15$ eV. The phase $\sigma_{1}$ is taken in such a way that it satisfies the relations given in Eq.(2.19). The viable values of the parameters $z_{1}$ and $z_{2}$ are calculated as functions of two mass squared differences and phase $\beta$, from the expressions equations (2.22) and (2.23) by making sure that it is consistent with the bound on total neutrino mass $\Sigma=m_{1}+m_{2}+m_{3}<0.12$ eV through the expression in Eq.(2.24). It is observed that CP asymmetries of the order $\mathcal{O}(10^{-7})$ can be produced for triplet masses $\sim 10^{10}$ GeV. The figures (8) and (9) are generated by taking $M_{1}$ ($M_{2}$) $=2\times 10^{10}$ GeV and taking the mass squared differences in $3\sigma$ range. The scale is suitable for studying two-flavor leptogenesis which we take up in the upcoming sections.
4.2 Baryogenesis from $\Delta_{1}$ decay
First, we consider the case -I, $M_{1}\ll M_{2}$, where the lepton asymmetry is generated by the decay processes of triplet $\Delta_{1}$. We obtain the final baryon asymmetry given in Eq. (3.24) via lepton asymmetries $\Delta_{\alpha}$ by solving the set of DMEs given in equations (3.16 - 3.19), for different sets of triplet mass $M_{1}$ and the flavored CP asymmetries $\epsilon^{\alpha\beta}_{1}$. The results of baryon asymmetry for some of the benchmark points are briefed in table (3). In table (3), we show the interplay of hierarchical branching ratios and the SCPV phase $\beta$ on the flavored CP asymmetry parameters. For example, for $M_{1}=3\times 10^{10}$ GeV, using Eq. (3.2), we calculate the branching ratios, $B_{1}^{L}$ using the allowed values of the parameters of the model. Using the same values of the parameters we also calculate the corresponding CP asymmetries as shown in the table. While the hierarchy in $B_{1}^{L}$ increases, it should have resulted in suppression of the magnitude of the CP asymmetries (as can be seen in Eq. (3.8)). But the contribution of the SCPV phase $\beta$ in $\epsilon^{\alpha\beta}_{1}$ (equations (C.8 - C.11)) compensates the suppression arising from the increase in the hierarchy of the branching ratios. It should be noted here that $B_{1}^{L}$ is independent of the phase $\beta$. This feature also enhances the efficiency of leptogenesis which can be seen from the increase in the final baryon asymmetry as shown in the table.
$M_{1}({\rm GeV})$
$B^{L}_{1}$
$\beta(^{\circ})$
$\epsilon^{ee}_{1}/10^{-8}$
$\epsilon^{e\mu}_{1}/10^{-9}$
$\epsilon^{\mu\mu}_{1}/10^{-9}$
$\epsilon^{\tau\tau}_{1}/10^{-9}$
$|\eta_{B}|/10^{-10}$
$0.48$
$178$
$0.70$
$0.84$
$1.01$
$-1.68$
$0.04$
$3\times 10^{10}$
$0.06$
$175$
$2.06$
$2.46$
$2.94$
$-4.91$
$0.29$
$0.005$
$165$
$2.04$
$2.43$
$2.91$
$-4.86$
$0.73$
$0.0005$
$126$
$2.06$
$2.46$
$2.94$
$-4.91$
$1.74$
$0.34$
$178$
$2.79$
$3.32$
$3.98$
$-6.63$
$0.30$
$4\times 10^{10}$
$0.05$
$175$
$2.82$
$3.36$
$4.03$
$-6.71$
$0.44$
$0.005$
$164$
$2.82$
$3.36$
$4.03$
$-6.71$
$1.07$
$0.0006$
$126$
$2.83$
$3.37$
$4.04$
$-6.74$
$2.50$
Table 3: Baryon asymmetries through flavored leptogenesis for the triplet mass hierarchy case-1, $M_{1}\ll M_{2}$ by solving DME in the energy range $[10^{9},10^{12}]$ GeV for different values of $M_{1}\sim 10^{10}$ GeV, $\epsilon^{\alpha\beta}_{1}$, and hierarchical $(B^{L}_{1}\sim 0.0005,0.005,0.05)$ as well as comparable $(B^{L}_{1}\sim 0.5)$ branching ratio $B^{L}_{1}$.
We present a specific result corresponding to a set with $M_{1}=3\times 10^{10}$ GeV, and flavored CP asymmetries $\epsilon^{ee}_{1}\sim 10^{-7}$, $\epsilon^{e\mu,~{}\mu e,~{}\mu\mu,~{}\tau\tau}_{1}\sim 10^{-8}$ as given in table (4). We have plotted the obtained flavored lepton asymmetries and baryon asymmetry in Fig. (3(a)) along with the evolution of different washout terms in Fig. (3(b)). Although we have shown the results for final baryon asymmetry for a few benchmark points, we have taken different data sets corresponding to particular flavored CP asymmetries for $M_{1}\sim 10^{10}$ GeV and hierarchical branching ratios and we have plotted the obtained baryon asymmetry for each set in Fig. (5(a)).
$M_{1}({\rm GeV})$
$B^{L}_{1}$
$|u_{1}|(\rm eV)$
$\epsilon^{ee}_{1}/10^{-7}$
$\epsilon^{e\mu}_{1}/10^{-8}$
$\epsilon^{\mu\mu}_{1}/10^{-8}$
$\epsilon^{\tau\tau}_{1}/10^{-8}$
$|\eta_{B}|/10^{-10}$
$3\times 10^{10}$ Fig. (4)
$0.58$
$4.80$
$4.47$
$5.32$
$6.38$
$-1.06$
$5.29$
Table 4: Baryon asymmetries through leptogenesis by solving DME in the energy range $[10^{9},10^{12}]$ GeV for $M_{1}=3\times 10^{10}$ GeV.
Figure 4: Solution of DMEs for asymmetries and washouts for set with $M_{1}=3\times 10^{10}$ GeV by using the flavored CP asymmetries from table (4). Left: The color lines represent different asymmetries. Using the numerically obtained asymmetries the final baryon asymmetry is calculated. The green lines correspond to the observational value of the BAU. Right: The color lines represent numerically obtained amount of washouts along different flavor directions in units of $sHz$.
4.3 Baryogenesis from $\Delta_{2}$ decay
We further consider the triplet mass hierarchy case (II), with $M_{2}\ll M_{1}$ and take different sets with triplet $\Delta_{2}$ mass $M_{2}\sim 10^{10}$ GeV, flavored CP asymmetries and hierarchical branching ratios. The results are similar to case (I). The results are briefed in table (5). In this case, for comparatively lower values of triplet mass, the model can result in adequate baryon asymmetry. We present a specific result corresponding to a set with $M_{2}=1.4\times 10^{10}$ GeV, and flavored CP asymmetries $\epsilon^{ee}_{2}\sim 10^{-7}$, $\epsilon^{e\mu,~{}\mu e,~{}\mu\mu}_{2}\sim 10^{-8}$, $\epsilon^{\tau\tau}_{2}\sim 10^{-6}$ as given in table (6). We have plotted the obtained flavored lepton asymmetries and baryon asymmetry in Fig. (4(a)) along with the evolution of different washout terms in Fig. (4(b)). Similar to the case -I, we have taken different data sets corresponding to particular flavored CP asymmetries for $M_{2}\sim 10^{10}$ GeV and hierarchical branching ratios and we have plotted the obtained baryon asymmetry for each set in Fig. (5(b)).
Figure 5: Solution of DMEs for asymmetries and washouts for set with $M_{2}=1.4\times 10^{10}$ GeV by using the flavored CP asymmetries from table (6). Left: The color lines represent different asymmetries. Using the numerically obtained asymmetries the final baryon asymmetry is calculated. The green lines correspond to the observational value of the BAU. Right: The color lines represent numerically obtained amount of washouts along different flavor directions in units of $sHz$.
$M_{2}({\rm GeV})$
$B^{L}_{2}$
$\beta(^{\circ})$
$\epsilon^{ee}_{2}/10^{-8}$
$\epsilon^{e\mu}_{2}/10^{-9}$
$\epsilon^{\mu\mu}_{2}/10^{-9}$
$\epsilon^{\tau\tau}_{2}/10^{-7}$
$|\eta_{B}|/10^{-10}$
$0.59$
$176$
$-0.66$
$-0.79$
$-0.95$
$0.75$
$0.04$
$1\times 10^{10}$
$0.06$
$172$
$-0.67$
$-0.80$
$-0.96$
$0.76$
$0.07$
$0.005$
$157$
$-0.67$
$-0.80$
$-0.96$
$0.75$
$0.86$
$0.0006$
$126$
$-0.73$
$-0.87$
$-1.05$
$0.75$
$3.26$
$0.53$
$177$
$-1.29$
$-1.53$
$-1.84$
$1.45$
$0.01$
$2\times 10^{10}$
$0.05$
$176$
$-1.42$
$-1.68$
$-2.02$
$1.59$
$0.03$
$0.006$
$156$
$-1.42$
$-1.69$
$-2.03$
$1.58$
$1.66$
$0.0005$
$126$
$-1.42$
$-1.69$
$-2.03$
$1.45$
$7.51$
Table 5: Baryon asymmetries through flavored leptogenesis for the triplet mass hierarchy case - II, $M_{2}\ll M_{1}$ by solving DME in the energy range $[10^{9},10^{12}]$ GeV for different values of $M_{2}\sim 10^{10}$ GeV, $\epsilon^{\alpha\beta}_{2}$, and hierarchical $(B^{L}_{2}\sim 0.0005,0.005,0.05)$ as well as comparable $(B^{L}_{2}\sim 0.5)$ branching ratio $B^{L}_{2}$.
$M_{2}({\rm GeV})$
$B^{L}_{2}$
$|u_{2}|(\rm eV)$
$\epsilon^{ee}_{2}/10^{-8}$
$\epsilon^{e\mu}_{2}/10^{-9}$
$\epsilon^{\mu\mu}_{2}/10^{-9}$
$\epsilon^{\tau\tau}_{2}/10^{-7}$
$|\eta_{B}|/10^{-10}$
$1.4\times 10^{10}$ Fig. (5)
$0.0006$
$48.1$
$-1.00$
$-1.20$
$-1.44$
$1.04$
$4.92$
Table 6: Baryon asymmetries through leptogenesis by solving DME in the energy range $[10^{9},10^{12}]$ GeV for $M_{2}=1.4\times 10^{10}$ GeV.
Figure 6: Plot of obtained baryon asymmetries $|\eta_{B}|$ for different masses of the triplet scalars in left (right) for $M_{1}(M_{2})$ in GeV. The figures also show a comparative study of the efficiency of leptogenesis and its dependence on the branching ratio hierarchy. The bound in green shows the range of the baryon asymmetry from observations.
4.4 Summary
The results obtained in this section can be summarized as follows:
  1. 1.
    The SCPV phase $\beta$ acts as a common link between the high energy CPV in leptogenesis and low energy effect such as neutrinoless double beta decay as shown in Fig. (0(b)).
  2. 2.
    The hierarchical branching ratios result in different amounts of washout effects along different flavor directions. The amount of washouts greatly influences the efficiency of leptogenesis. Comparing figures (3(b)) ($B_{1}^{L}=0.58$) and (4(b)) ($B_{2}^{L}=0.0006$), one can see that in the first case, the washout effects along the flavor directions are more as compared to the second case (except $ee$-flavor direction). This feature can be attributed to the specific flavor structure of the triplet Yukawa coupling matrices of the model. Although, in the second case, the extreme hierarchy in the branching ratios can in principle suppress the flavored CP asymmetry parameters, the presence of SCPV phase $\beta$ can neutralise the effect.
  3. 3.
    The process of generation of baryon asymmetry from lepton asymmetry requires that CP violation should be maximal. Observing Eq. (3.8), it is ideal to generate maximal CP asymmetry for lower values of $M_{1}(M_{2})$. This would require $B^{L}\sim B^{\phi}$. In our analysis, in bringing down the $M_{1}(M_{2})$ scale as low as possible it can be observed that
    1. (a)
      In case - I, $M_{1}\ll M_{2}$, (in Fig. (4), also Fig. (5(a))) the observable range of baryon asymmetry can be achieved for triplet mass $M_{1}$ as low as $\sim 3\times 10^{10}$ GeV for comparable branching ratios $B^{L}_{1}\sim 0.5$. Taking hierarchical branching ratios and requiring adequate flavored CP asymmetries, baryon asymmetry can only be produced in the observable range for a triplet mass $M_{1}>5\times 10^{10}$ GeV for a highly hierarchical branching ratio $B^{L}_{1}\sim 0.0005$.
    2. (b)
      In case - II, $M_{2}\ll M_{1}$, we find from Fig. 5 ( also Fig. (5(b))) the observational range of baryon asymmetry can be achieved for triplet mass $M_{2}$ as low as $\sim 1\times 10^{10}$ GeV for highly hierarchical branching ratio $B^{L}_{2}\sim 0.0005$.
    3. (c)
      From Fig. (6), for both cases - I and II, with hierarchical branching ratios the baryon asymmetry increases for a fixed value of $M_{1}(M_{2})$. As shown in tables (3) and (5), the suppression in CP asymmetry due to the hierarchical branching ratio, can be compensated by the extra source of CP violation due to the presence of the SCPV phase $\beta$ in the neutrino mass matrix. As a result, even with the modified washout effect from hierarchical branching ratios, it is possible to achieve maximal CP violation and baryon asymmetry can be achieved in the observable range. The flavor structure of the model is suitable to draw these conclusions.
5 Conclusion
Spontaneous violation of CP symmetry is very promising in explaining CP violation in low-energy processes like neutrino oscillation and high-energy process like leptogenesis. In general CP violation in these two sectors are not related [21] and are model-dependent. In this paper, we have studied, in an extended Standard model, neutrino mass generated via type-II seesaw mechanism and leptogenesis from the decay of heavy scalar triplet with a common source of CP violation. The model provides a viable explanation of baryon asymmetry through thermal leptogenesis, by sourcing enough CP asymmetry spontaneously broken due to the presence of a scalar singlet with complex VEV. Moreover, the neutrino mass matrix exhibits TBM structure which generally fails to address the occurrence of non-zero CHOOZ mixing angle, $\theta_{13}$. By incorporating small perturbations around flavon VEV in the high-energy scale, it can accommodate $\theta_{13}\neq 0$ in the model. With the latest neutrino oscillation data, bound on total neutrino mass and the effective neutrino mass, we have analyzed the parameter space of the model. Using the constrained parameter space of the model we have studied flavored leptogenesis from the decay of the triplet scalars. The flavor structure of the model suppresses flavor-independent CP asymmetry and hence we explore the effect of flavor in leptogenesis by using a set of flavor-covariant Boltzmann equations considering a density matrix formalism. In our study, the scale of leptogenesis is $\mathcal{O}(10^{10})$ GeV. We also study the role of hierarchical branching ratios and the SCPV phase $\beta$ in generating maximal CP asymmetry to generate baryon asymmetry in the observational range. Requiring mass of triplet as low as possible, we find BAU in the observable range can be produced for $M_{1}=3\times 10^{10}$ GeV for comparable branching ratios. Whereas the same can be achieved for $M_{2}=1.4\times 10^{10}$ GeV for hierarchical branching ratios. The flavor structure of the triplet Yukawa couplings plays an important role in generating the washout effects which are inevitable for thermal leptogenesis.
Appendix A Scalar potential of the model
The scalar potential invariant under the symmetries as given in table (1), can be generally realised as
$V=V_{S}+V_{\phi}+V_{\Delta}+V_{S\phi}+V_{S\Delta}+V_{\phi\Delta}+V_{S\phi\Delta},$ (A.1)
where
$V_{S}$ $=\mu^{2}_{S}(S^{2}+S^{*2})+m^{2}_{S}S^{*}S+\lambda^{\prime}_{S}(S^{4}+S^{*4})+\lambda^{\prime\prime}_{S}S^{*}S(S^{2}+S^{*2})+\lambda_{S}(S^{*}S)^{2},$ (A.2)
$V_{\phi}$ $=m^{2}_{\phi}\phi^{\dagger}\phi+\lambda_{\phi}(\phi^{\dagger}\phi)^{2},$ (A.3)
$V_{\Delta}$ $=\Sigma_{a}M^{2}_{a}Tr(\Delta^{\dagger}_{a}\Delta_{a})+\Sigma_{a,b}[(\lambda_{\Delta})_{ab}Tr(\Delta^{\dagger}_{a}\Delta_{a})Tr(\Delta^{\dagger}_{b}\Delta_{b})+(\lambda^{\prime}_{\Delta})_{ab}Tr(\Delta^{\dagger}_{a}\Delta_{a}\Delta^{\dagger}_{b}\Delta_{b})],$ (A.4)
$V_{S\phi}$ $=\eta_{S}(S^{*}S)(\phi^{\dagger}\phi)+\eta^{\prime}_{S}(S^{2}+S^{*2})(\phi^{\dagger}\phi),$ (A.5)
$V_{S\Delta}$ $=\Sigma_{a}Tr(\Delta^{\dagger}_{a}\Delta_{a})[\eta_{a}(S{2}+S^{*2})+\xi_{a}S^{*}S],$ (A.6)
$V_{\phi\Delta}$ $=\Sigma_{a}[\xi^{\prime}_{a}(\phi^{\dagger}\phi)Tr(\Delta^{\dagger}_{a}\Delta_{a})+\xi^{\prime\prime}_{a}(\phi^{\dagger}\Delta^{\dagger}_{a}\Delta_{a}\phi)]+(\mu_{2}M_{2}\tilde{\phi}^{T}\Delta_{2}\tilde{\phi}+{\rm H.c.}),$ (A.7)
$V_{S\phi\Delta}$ $=\tilde{\phi}^{T}\Delta_{1}\tilde{\phi}(\lambda_{1}S+\lambda^{\prime}_{1}S^{*})+{\rm H.c.},$ (A.8)
where $\tilde{\phi}=i\sigma_{2}\phi^{*}$. It is interesting to note that a viable leptogenesis scenario in this context requires a term $\mu_{1}M_{1}\tilde{\phi}^{T}\Delta_{1}\tilde{\phi}$ such a term is forbidden due to the choice of $Z_{4}$ symmetry. However, the complex VEV acquired by the scalar singlet $S$ enables this term to be generated from the scalar potential as shown in Eq. (A.8). Furthermore, the addition of two heavy scalar fields allows us to explore the spontaneous breaking of the discrete $A_{4}$ symmetry.
The potential can be written in tree level as
$V_{0}=m^{2}_{S}v^{2}_{S}+\lambda_{S}v^{4}_{S}+2(\mu^{2}_{S}+\lambda^{\prime\prime}_{S}v^{2}_{S})v^{2}_{S}\cos(2\alpha)+2\lambda^{\prime}_{S}v^{4}_{S}\cos(4\alpha).$ (A.9)
Minimizing the potential with respect to $v_{S}$ and $\alpha$ we obtain,
$\frac{\partial V_{0}}{\partial v_{S}}=2v_{S}[m^{2}_{S}+2\lambda_{S}v^{2}_{S}+2(\mu^{2}_{S}+2\lambda^{\prime\prime}_{S}v^{2}_{S})\cos(2\alpha)+4\lambda^{\prime}_{S}v^{2}_{S}\cos(4\alpha)]=0,$ (A.10)
and
$\frac{\partial V_{0}}{\partial\alpha}=-4v^{2}_{S}\sin(2\alpha)[(\mu^{2}_{S}+\lambda^{\prime\prime}_{S}v^{2}_{S})+4\lambda^{\prime}_{S}v^{2}_{S}\cos(2\alpha)]=0.$ (A.11)
Three possible solution to the above equations, excluding the trivial solution $v_{S}=0$ that leads to $V_{0}=0$, are given by
  1. 1.
    $v^{2}_{S}=-\frac{m^{2}_{S}+2\mu^{2}_{S}}{2(\lambda_{S}+2\lambda^{\prime}_{S}+2\lambda^{\prime\prime}_{S})},\quad\alpha=0,\pm\pi,$ (A.12)
  2. 2.
    $v^{2}_{S}=\frac{-m^{2}_{S}+2\mu^{2}_{S}}{2(\lambda_{S}+2\lambda^{\prime}_{S}-2\lambda^{\prime\prime}_{S})},\quad\alpha=\pm\frac{\pi}{2},$ (A.13)
  3. 3.
    $v^{2}_{S}=\frac{-2\lambda^{\prime}_{S}m^{2}_{S}+\lambda^{\prime\prime}_{S}\mu^{2}_{S}}{4\lambda_{S}\lambda^{\prime}_{S}-8\lambda^{\prime}_{S}-\lambda^{\prime\prime}_{S}},\quad\cos(2\alpha)=-\frac{\mu^{2}_{S}+\lambda^{\prime\prime}_{S}v^{2}_{S}}{4\lambda^{\prime}_{S}v^{2}_{S}}.$ (A.14)
To show that the last solution, which also leads in this case to the spontaneous breaking of the CP symmetry, corresponds to the global minimum of the potential, we consider $m^{2}_{S}<0$, $\lambda^{\prime\prime}_{S}\simeq 0$, $\mu_{S}\simeq 0$ and $\lambda_{S}>2\lambda^{\prime}_{S}>0$ and we obtain from Eq. (A.13),
$v^{2}_{S}\simeq-\frac{m^{2}_{S}}{2(\lambda_{S}+2\lambda^{\prime}_{S})},\quad\alpha=0,\pm\frac{\pi}{2},\pm\pi.$ (A.15)
For cases - 1, and 2,
$V_{0}\simeq-\frac{m^{4}_{S}}{4(\lambda_{S}+2\lambda^{\prime}_{S})}.$ (A.16)
For case - 3
$v^{2}_{S}\simeq-\frac{m^{2}_{S}}{2(\lambda_{S}-2\lambda^{\prime}_{S})},\quad\alpha\simeq\pm\frac{\pi}{4},$ (A.17)
which leads to
$V_{0}\simeq-\frac{m^{4}_{S}}{4(\lambda_{S}-2\lambda^{\prime}_{S})}$ (A.18)
which is the absolute minimum of the potential.
Appendix B Reaction Densities
The space-time density of total decay of triplet scalar $\Delta$ and its anti-particle $\overline{\Delta}$, $\gamma_{D}$ and that of $2\leftrightarrow 2$ scattering, $\gamma_{A}$ are given by,
$\gamma_{D}=s\Gamma_{\Delta}\Sigma^{eq}_{\Delta}\frac{K_{1}(z)}{K_{2}(z)},$ (B.1)
[…]
Here, $x={\rm s}/M^{2}_{a}$, s is the Mandelstam variable, square of center of mass energy [29], $\hat{\sigma}_{S}$ is the reduced cross-section of $2\rightarrow 2$ scatterings (relevant diagrams in fig.(7)). For processes mediated by gauge bosons $x_{\rm min}=4$ and for Yukawa-induced reactions $x_{\rm min}=0$.
The corresponding reaction density of scattering is given by,
$\gamma_{A}=\frac{M_{a}T^{3}e^{-\frac{2M_{a}}{T}}}{64\pi^{4}}(9g^{4}+12g_{2}^{2}g^{2}_{\mathcal{Y}}+3g^{4}_{Y})\left(1+\frac{3T}{4M_{a}}\right).$ (B.3)
Figure 7: Feynman diagrams that contribute to the interaction rate $\gamma_{A}$
Here we take $\Gamma_{\Delta}=\frac{M^{2}_{a}\tilde{m}_{\Delta}}{16\pi v^{2}\sqrt{B^{L}B^{\phi}}}$, $\tilde{m}_{\Delta}={\rm Tr}(m^{(a)\dagger}_{\nu}m^{(a)}_{\nu})$, $\Sigma^{\rm eq}_{\Delta}=2Y^{\rm eq}_{\Delta}=2\times\frac{45g_{\Delta}}{4\pi^{4}g_{*}}z^{2}K_{2}(z)$ and $g_{\Delta}=3$, $g_{*}=106.75$ are the triplet scalar and SM particle-degrees of freedom, respectively [26]. $g_{2},g_{\mathcal{Y}}$ are the coupling constants corresponding to $SU(2)_{L}$ and $U(1)_{\mathcal{Y}}$ interactions, respectively, having values, $g_{2}=\frac{e}{\sin\theta_{W}}=0.651742$ and $g_{\mathcal{Y}}=\frac{e}{\cos\theta_{W}}=0.461388$. The other reaction densities, $\gamma$’s are to be calculated using the integration given in Eq. (B.2).
Within the temperature range $10^{9}$ GeV $<T<10^{12}$ GeV, the chemical potential relations are given by:
$\sum_{i=1,2,3}(\mu_{q_{i}}+2\mu_{u_{i}}-\mu_{d_{i}})-\sum_{\alpha=e,\mu,\tau}(\mu_{l_{\alpha}}+\mu_{e_{\alpha}})+2\mu_{\phi}6\mu_{\Delta}=0\quad(\Delta_{Y}=0)$ (B.4)
$\sum_{i=1,2,3}(2\mu_{q_{i}}+\mu_{u_{i}}+\mu_{d_{i}})=0\quad(\Delta_{B}=0)$ (B.5)
$\sum_{i=1,2,3}(2\mu_{q_{i}}-\mu_{u_{i}}-\mu_{d_{i}})=0\quad(\text{QCD sphaleron})$ (B.6)
$\sum_{i=1,2,3}2\mu_{q_{i}}+\sum_{\alpha=e,\mu,\tau}\mu_{l_{\alpha}}=0\quad(\text{EW sphaleron})$ (B.7)
[…]
$\mu_{l_{\alpha}}-\mu_{e_{\alpha}}-\mu_{\phi}=0\quad(\text{charged lepton Yukawa})$ (B.10)
Appendix C Estimation of flavored CP asymmetries
The important factors appearing in the flavored CP asymmetry parameters as given in Eq. (3.8) are calculated as follows:
$\rm{Im}\left[m^{(1)}_{\nu,ee}m^{*}_{\nu,ee}\right]=\frac{2}{3}z_{1}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\left(1-\varepsilon_{1}-\varepsilon_{2}\right)\sin(\beta-\sigma_{1}),$ (C.1)
$\rm{Im}\left[m^{(1)}_{\nu,e\mu}m^{*}_{\nu,e\mu}\right]$ $=$ $-\frac{2}{3}\varepsilon_{1}z_{1}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})$ (C.2)
$=$ $\rm{Im}\left[m^{(1)}_{\nu,\mu e}m^{*}_{\nu,\mu e}\right],$
$\rm{Im}\left[m^{(1)}_{\nu,\mu\mu}m^{*}_{\nu,\mu\mu}\right]$ $=$ $\varepsilon_{2}z_{1}\Big{[}\frac{1}{3}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})$ (C.3)
$+$ $\left(z^{2}_{1}+z^{2}_{2}-2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{2})\Big{]},$
$\rm{Im}\left[m^{(1)}_{\nu,\tau\tau}m^{*}_{\nu,\tau\tau}\right]$ $=$ $\varepsilon_{1}z_{1}\Big{[}\frac{1}{3}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})$ (C.4)
$+$ $\left(z^{2}_{1}+z^{2}_{2}-2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{2})\Big{]},$
$\rm{Im}\left[m^{(2)}_{\nu,ee}m^{*}_{\nu,ee}\right]$ $=$ $-\frac{2}{3}z_{2}\Big{[}\frac{2}{3}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}(1-2\varepsilon_{1}-2\varepsilon_{2})\sin\sigma_{1}$ (C.5)
$+$ $\frac{1}{3}z_{1}(1+\varepsilon_{1}+\varepsilon_{2})\sin\beta\Big{]},$
$\rm{Im}\left[m^{(2)}_{\nu,e\mu}m^{*}_{\nu,e\mu}\right]$ $=$ $\frac{1}{3}z_{2}\Big{[}(\frac{7}{6}\varepsilon_{1}+\frac{1}{6}\varepsilon_{2}-\frac{1}{3})\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin\sigma_{1}$
$+$ $\frac{1}{3}(1+\varepsilon_{1}+\varepsilon_{2})z_{1}\sin\beta$
-- $\frac{1}{2}(\varepsilon_{1}-\varepsilon_{2})\left(z^{2}_{1}+z^{2}_{2}-2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin\sigma_{2}\Big{]}$
$=$ $\rm{Im}\left[m^{(2)}_{\nu,\mu e}m^{*}_{\nu,\mu e}\right],$
$\rm{Im}\left[m^{(2)}_{\nu,\mu\mu}m^{*}_{\nu,\mu\mu}\right]$ $=$ $-\frac{2}{3}z_{2}\Big{[}(-\frac{5}{6}\varepsilon_{1}+\frac{1}{6}\varepsilon_{2}-\frac{1}{6})\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin\sigma_{1}$ (C.6)
$+$ $\frac{1}{3}(1+\varepsilon_{1}+\varepsilon_{2})z_{1}\sin\beta$
$+$ $\frac{1}{2}(1-\varepsilon_{1}-3\varepsilon_{2})\left(z^{2}_{1}+z^{2}_{2}-2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin\sigma_{2}\Big{]},$
and
$\rm{Im}\left[m^{(2)}_{\nu,\tau\tau}m^{*}_{\nu,\tau\tau}\right]$ $=$ $-\frac{2}{3}z_{2}\Big{[}(\frac{1}{6}\varepsilon_{1}-\frac{5}{6}\varepsilon_{2}+\frac{1}{6})\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin\sigma_{1}$ (C.7)
$+$ $\frac{1}{3}(1+\varepsilon_{1}+\varepsilon_{2})z_{1}\sin\beta$
$+$ $\frac{1}{2}(1-3\varepsilon_{1}-\varepsilon_{2})\left(z^{2}_{1}+z^{2}_{2}-2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin\sigma_{2}\Big{]}.$
Finally from equations (4.1), (C.1), (C.2), (C.3) and (C.4), the flavored CP asymmetries $\epsilon^{\alpha\beta}_{1}$ are formed as,
$\epsilon^{ee}_{1}=\frac{M^{2}_{1}|u_{1}|^{2}}{2\pi\left(3z^{2}_{1}v^{4}+4M^{4}_{1}|u_{1}|^{4}\right)}\Big{[}\frac{2}{3}z_{1}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\left(1-\varepsilon_{1}-\varepsilon_{2}\right)\sin(\beta-\sigma_{1})\Big{]},$ (C.8)
$\epsilon^{e\mu}_{1}$$\approx$ $\frac{M^{2}_{1}|u_{1}|^{2}}{2\pi\left(3z^{2}_{1}v^{4}+4M^{4}_{1}|u_{1}|^{4}\right)}\Big{[}-\frac{2}{3}\varepsilon_{1}z_{1}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})\Big{]}$ (C.9)
$\approx$ $\epsilon^{\mu e}_{1},$
$\epsilon^{\mu\mu}_{1}$$\approx$ $\frac{M^{2}_{1}|u_{1}|^{2}}{2\pi\left(3z^{2}_{1}v^{4}+4M^{4}_{1}|u_{1}|^{4}\right)}\times\varepsilon_{2}z_{1}\Big{[}\frac{1}{3}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})$ (C.10)
$+$ $\left(z^{2}_{1}+z^{2}_{2}-2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{2})\Big{]},$
and
$\epsilon^{\tau\tau}_{1}$$\approx$ $\frac{M^{2}_{1}|u_{1}|^{2}}{2\pi\left(3z^{2}_{1}v^{4}+4M^{4}_{1}|u_{1}|^{4}\right)}\times\varepsilon_{1}z_{1}\Big{[}\frac{1}{3}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})$ (C.11)
$+$ $\left(z^{2}_{1}+z^{2}_{2}-2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{2})\Big{]}.$
From equations (4.2), (C.5), (C), (C.6) and (C.7), the flavored CP asymmetries $\epsilon^{\alpha\beta}_{2}$ are formed as,
$\epsilon^{ee}_{2}$ $=$ $\frac{M^{2}_{2}|u_{2}|^{2}}{4\pi\left(z^{2}_{2}v^{4}(1-2\varepsilon_{1}-2\varepsilon_{2})+2M^{4}_{2}|u_{2}|^{4}\right)}\times-\frac{2}{3}z_{2}$ (C.12)
$\Big{[}\frac{2}{3}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}(1-2\varepsilon_{1}-2\varepsilon_{2})\sin\sigma_{1}$
$+$ $\frac{1}{3}z_{1}(1+\varepsilon_{1}+\varepsilon_{2})\sin\beta\Big{]},$
$\epsilon^{e\mu}_{2}$$\approx$ $\frac{M^{2}_{2}|u_{2}|^{2}}{4\pi\left(z^{2}_{2}v^{4}(1-2\varepsilon_{1}-2\varepsilon_{2})+2M^{4}_{2}|u_{2}|^{4}\right)}\times\frac{1}{3}z_{2}$ (C.13)
$\Big{[}(\frac{7}{6}\varepsilon_{1}+\frac{1}{6}\varepsilon_{2}-\frac{1}{3})\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin\sigma_{1}$
$+$ $\frac{1}{3}(1+\varepsilon_{1}+\varepsilon_{2})z_{1}\sin\beta$
-- $\frac{1}{2}(\varepsilon_{1}-\varepsilon_{2})\left(z^{2}_{1}+z^{2}_{2}-2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin\sigma_{2}\Big{]}$
$\approx$ $\epsilon^{\mu e}_{2},$
$\epsilon^{\mu\mu}_{2}$$\approx$ $\frac{M^{2}_{2}|u_{2}|^{2}}{4\pi\left(z^{2}_{2}v^{4}(1-2\varepsilon_{1}-2\varepsilon_{2})+2M^{4}_{2}|u_{2}|^{4}\right)}\times-\frac{2}{3}z_{2}$ (C.14)
$\Big{[}(-\frac{5}{6}\varepsilon_{1}+\frac{1}{6}\varepsilon_{2}-\frac{1}{6})\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin\sigma_{1}$
$+$ $\frac{1}{3}(1+\varepsilon_{1}+\varepsilon_{2})z_{1}\sin\beta$
$+$ $\frac{1}{2}(1-\varepsilon_{1}-3\varepsilon_{2})\left(z^{2}_{1}+z^{2}_{2}-2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin\sigma_{2}\Big{]},$
and
$\epsilon^{\tau\tau}_{2}$$\approx$ $\frac{M^{2}_{2}|u_{2}|^{2}}{4\pi\left(z^{2}_{2}v^{4}(1-2\varepsilon_{1}-2\varepsilon_{2})+2M^{4}_{2}|u_{2}|^{4}\right)}\times-\frac{2}{3}z_{2}$ (C.15)
$\Big{[}(\frac{1}{6}\varepsilon_{1}-\frac{5}{6}\varepsilon_{2}+\frac{1}{6})\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin\sigma_{1}$
$+$ $\frac{1}{3}(1+\varepsilon_{1}+\varepsilon_{2})z_{1}\sin\beta$
$+$ $\frac{1}{2}(1-3\varepsilon_{1}-\varepsilon_{2})\left(z^{2}_{1}+z^{2}_{2}-2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin\sigma_{2}\Big{]}.$
(a) $\epsilon^{ee}_{1}$ as a function of $\beta$
(b) $\epsilon^{e\mu}_{1}$ as a function of $\beta$
(c) $\epsilon^{\mu\mu}_{1}$ as a function of $\beta$
(d) $\epsilon^{\tau\tau}_{1}$ as a function of $\beta$
Figure 8: Flavored CP asymmetries generated due to triplet $\Delta_{1}$ decays, as a function of phase $\beta$ for comparable and hierarchical branching ratios. The triplet $\Delta_{1}$ mass $M_{1}$ is taken $2\times 10^{10}$ GeV.
(a) $\epsilon^{ee}_{2}$ as a function of $\beta$
(b) $\epsilon^{e\mu}_{2}$ as a function of $\beta$
(c) $\epsilon^{\mu\mu}_{2}$ as a function of $\beta$
(d) $\epsilon^{\tau\tau}_{2}$ as a function of $\beta$
Figure 9: Flavored CP asymmetries generated due to triplet $\Delta_{2}$ decays, as a function of phase $\beta$ for comparable and hierarchical branching ratios. The triplet $\Delta_{2}$ mass $M_{2}$ is taken $2\times 10^{10}$ GeV.
Acknowledgments
SC acknowledges the valuable discussions with Rohan Pramanick on Spontaneous CP violation and leptogenesis.

### v2

@@
 [31] that even for extremely hierarchical branching ratios, it is possible to achieve the required BAU.
-In this work, we focus on leptogenesis from the decay of triplet scalars with hierarchical branching ratios and CP violation originating from the VEV of a scalar singlet, embedded in the SM along with with two triplet scalars.
+In this work, we focus on leptogenesis from the decay of triplet scalars with hierarchical branching ratios and CP violation originating from the VEV of a scalar singlet, embedded in the SM along with two triplet scalars.
 Also to make the model more predictive we impose $A_{4}$ symmetry at the Lagrangian level.
@@
 The model is introduced in section (2).
-Section (3) is devoted to pointing out the important features of flavored CP asymmetry and DME for the analysis.
+Section (3) is devoted to point out the important features of flavored CP asymmetry and DME for the analysis.
 Finally, the results are presented in section (4) with the summary.
@@
 Three appendices (A), (B), (C) are added containing some of the detailed calculations.
-Field
-$L$
-$e_{R},\mu_{R},\tau_{R}$
-$\Delta_{1}$
-$\Delta_{2}$
-$\phi$
-$S$
-$\Phi$
-$\psi$
-$SU(2)_{L}$
-$2$
-$1$
-$3$
-$3$
-$2$
-$1$
-$1$
-$1$
-$U(1)_{Y}$
-$-1/2$
-$-1$
-$1$
-$1$
-$1/2$
-00
-00
-00
-$A_{4}$
-$3$
-$1,1^{\prime\prime},1^{\prime}$
-$1$
-$1$
-$1$
-$1$
-$3$
-$3$
-$Z_{4}$
-$i$
-$-i$
-$1$
-$-1$
-$i$
-$-1$
-$i$
-$1$
+Field $L$ $e_{R},\mu_{R},\tau_{R}$ $\Delta_{1}$ $\Delta_{2}$ $\phi$ $S$ $\Phi$ $\psi$
+$SU(2)_{L}$ $2$ $1$ $3$ $3$ $2$ $1$ $1$ $1$
+$U(1)_{Y}$ $-1/2$ $-1$ $1$ $1$ $1/2$ 00 00 00
+$A_{4}$ $3$ $1,1^{\prime\prime},1^{\prime}$ $1$ $1$ $1$ $1$ $3$ $3$
+$Z_{4}$ $i$ $-i$ $1$ $-1$ $i$ $-1$ $i$ $1$
 Table 1: Field representations under $SU(2)_{L}$, $U(1)_{Y},A_{4}$ and $Z_{4}$ symmetries.
@@
 which leads to
-$V_{0}\simeq-\frac{m^{4}_{S}}{4(\lambda_{S}-2\lambda^{\prime}_{S})}$ (2.11)
+$V_{0}\simeq-\frac{m^{4}_{S}}{4(\lambda_{S}-2\lambda^{\prime}_{S})},$ (2.11)
 which is the absolute minimum of the potential.
+One of the constraints on models with spontaneous CP violation comes from the cosmological scenario of domain wall problem [54].
+A review of topological defects like cosmic strings, domain walls, and monopoles arising from cosmological phase transition in the early universe can be found in Ref.
+[53].
+The spontaneous breaking of a global discrete symmetry leads to the formation of domain walls [50, 51].
+The energy density of the domain walls redshifts slower than those of matter and radiation.
+Long-lived domain walls conflict with standard big bang cosmology.
+There are several solutions to make them unstable.
+One of the solutions is assuming gravity breaks global discrete symmetries explicitly, which lifts the degeneracy of the two vacua, there will be no domain wall problem [52, 49].
+Lifting the degeneracy of the two vacua by an amount $\epsilon$ generates a pressure difference between the two sides of the walls with a tendency to push the wall into the false vacuum.
+By making a careful analysis of the dynamics of evolution with domain walls in a radiation-dominated universe, it can be shown that the domain walls are pushed to false vacuum if
+$\epsilon\geq\frac{v_{\rm dis}^{6}}{M^{2}_{\rm Pl}},$ (2.12)
+where $v_{\rm dis}$ is the scale of discrete symmetry breaking and $M_{\rm Pl}\sim 10^{19}$ GeV.
+In our case, $v_{\rm dis}$ can be replaced by the SCPV scale, $v_{S}$.
+The authors of Ref.
+[52] have considered generating such gravity-induced terms and their effects in destabilizing the domain walls.
+For a theory with generic neutral scalar field $\phi$ the Planck-scale suppressed/gravity-induced operators can be written as
+$\Delta V=\frac{c_{5}}{M_{\rm Pl}}\phi^{5}+\frac{c_{6}}{M^{2}_{\rm Pl}}\phi^{6}+{\rm Higher~{}order~{}terms}.$ (2.13)
+In general, the coefficients $c_{n}$ are of order $1$.
+The terms on the right-hand side of Eq.
+(2.13) are generic.
+In a specific model such as the one studied in this paper, the effective higher degree potential induced by gravity has to be constructed carefully keeping gauge invariance intact.
+In the model under consideration, the Lagrangian has been written keeping $A_{4}\times Z_{4}$, CP and gauge invariance intact.
+Also, we have assumed that the singlet scalar $S$ is very heavy and decouples from the theory at an energy scale much higher than the electroweak and Higgs-triplet scales.
+In Appendix (A), the minimization of the scalar potential is done by taking only the complex vev of $S$ into account.
+Keeping these in mind, the first higher-dimensional operator that can break CP symmetry explicitly can be written in terms of the complex gauge singlet as
+$\Delta V=\frac{c_{6}}{M^{2}_{\rm Pl}}(SS^{*})^{3}+{\rm H.c.}+{\rm Higher~{}order~{}terms}.$ (2.14)
+Since gravity is expected to break discrete symmetry explicitly, CP symmetry breaking can be ensured by taking $c_{6}$ as complex.
+However, the successful removal of domain walls requires $c_{6}\gtrsim 1$ as per Eq.(2.12).
+So careful study of the detailed consequences of gravity is important in low energy issues such as CP violation.
+Nevertheless, gravity can break the $Z_{4}$ discrete symmetry ($S$ is charged under $Z_{4}$).
+Thus, in this case, it is possible to construct a dimension-five operator of the form
+$\Delta V=\frac{c_{5}}{M_{\rm Pl}}\mathcal{O}+{\rm H.c.}+{\rm Higher~{}order~{}terms},$ (2.15)
+where the operator $\mathcal{O}$ can be of the form $S^{5},(SS^{*})S^{3}$, and $(SS^{*})^{2}S$.
+CP symmetry is explicitly broken with the dimension-five operator with $c_{5}$ as complex.
+However by comparing the above contribution as per Eq.(2.12), the relation $c_{5}\gg\frac{v_{S}}{M_{\rm Pl}}$ has to hold.
+This ensures the dimension-five operator is much bigger than the amount $\frac{v_{S}^{6}}{M^{2}_{\rm Pl}}$.
+In our case the scale of leptogenesis is $>10^{10}$ GeV.
+So taking the $v_{S}$, the SCPV scale above the leptogenesis scale, the values of $c_{5}$ are bounded below as
+$c_{5}>\begin{cases}10^{-7},&v_{S}\sim 10^{12},\\ 10^{-6},&v_{S}\sim 10^{13},\\ 10^{-5},&v_{S}\sim 10^{14}.\end{cases}$ (2.16)
+The above bound is easily satisfied as $c_{5}$ is of order $1$. This makes the domain wall network disappear early enough for an acceptable leptogenesis scenario.
+There are several other ways to deal with the domain wall problem.
+The domain walls can become unstable by the existence of a non-zero $\theta_{\rm QCD}$ through non-perturbative effects, that break the CP explicitly and lift the degeneracy of the two CP conjugate vacua [55].
+Another idea in the same direction is to invoke inflation after the domain walls are produced [56].
+In Ref.
+[57], the authors study the possibility of spontaneous discrete symmetry breaking due to a gauge singlet field whose initial value is fixed during inflation, so that the resulting discrete symmetry-breaking phase is the same over the whole observable universe.
+However, the baryon asymmetry has to be generated after inflation to avoid diluting it to an insignificant level.
+In most of the models on leptogenesis, the reheat temperature required is $10^{12}$ GeV (in our case the scale of leptogenesis is $\mathcal{O}(10^{10})$ GeV).
+So it is hard to make domain walls unstable unless they are associated with a spontaneous breaking $>10^{12}$ GeV which fits very well in our case.
+The various ways to solve domain wall problem show that it is possible to overcome it independently.
+In all the scenarios considered above, it does not prevent the complex VEV of $S$ in generating a complex Yukawa coupling required for CP violation in leptogenesis.
 The triplet Higgs scalars contribute to the neutrino mass generation through the type-II seesaw mechanism as
@@
 where $m^{(1)}_{\nu}$ and $m^{(2)}_{\nu}$ are the contributions coming from the triplet scalars $\Delta_{1}$ and $\Delta_{2}$, respectively.
-Here, $u_{a}$ is the VEV of the neutral components of the triplet scalars $\Delta_{a}$, and it is given by
+Here, $u_{a}$ are the VEVs of the neutral components of the triplet scalars $\Delta_{a}$, and it is given by
 $u_{a}=\mu^{*}_{a}v^{2}/M_{a},\quad\mu_{1}=(\lambda_{1}e^{i\alpha}+\lambda^{\prime}_{1}e^{-i\alpha})v_{S}/M_{1}$ (2.18)
@@
 TBM mixing matrix has gathered considerable attention because of its proximity with the observation of neutrino masses and mixing.
-However the absence of Dirac-type CP violating phase asks for possible modifications of the model, which we discuss in the next subsection.
+However, the absence of Dirac-type CP violating phase asks for possible modifications of the model, which we discuss in the next subsection.
 Considering the two cases of neutrino mass hierarchy, we can have two sets of constraints on $z_{1}$ and $z_{2}$.
@@
 NuFIT 5.0 [24]
-Neutrino Oscillation parameters
 […]
-[…]
-$\sin^{2}\theta_{12}$
-$0.303^{+0.012}_{-0.012}$
-$0.270\rightarrow 0.341$
-$\sin^{2}\theta_{23}$
-$0.451^{+0.019}_{-0.016}$
-$0.408\rightarrow 0.603$
-$\sin^{2}\theta_{13}$
-$0.02225^{+0.00056}_{-0.00059}$
-$0.02052\rightarrow 0.02398$
-$\delta/^{\circ}$
-$232^{+36}_{-26}$
-$144\rightarrow 350$
-$\frac{\Delta m^{2}_{21}}{10^{-5}{\rm eV}^{2}}$
-$7.41^{+0.21}_{-0.20}$
-$6.82\rightarrow 8.03$
-$\frac{\Delta m^{2}_{3l}}{10^{-3}{\rm eV}^{2}}$
-$+2.507^{+0.026}_{-0.027}$
-$+2.427\rightarrow+2.590$
+$\sin^{2}\theta_{12}$ $0.303^{+0.012}_{-0.012}$ $0.270\rightarrow 0.341$
+$\sin^{2}\theta_{23}$ $0.451^{+0.019}_{-0.016}$ $0.408\rightarrow 0.603$
+$\sin^{2}\theta_{13}$ $0.02225^{+0.00056}_{-0.00059}$ $0.02052\rightarrow 0.02398$
+$\delta/^{\circ}$ $232^{+36}_{-26}$ $144\rightarrow 350$
+$\frac{\Delta m^{2}_{21}}{10^{-5}{\rm eV}^{2}}$ $7.41^{+0.21}_{-0.20}$ $6.82\rightarrow 8.03$
+$\frac{\Delta m^{2}_{3l}}{10^{-3}{\rm eV}^{2}}$ $+2.507^{+0.026}_{-0.027}$ $+2.427\rightarrow+2.590$
 Table 2: Allowed ranges of neutrino oscillation parameters for normal mass ordering.
@@
 By imposing the experimentally obtained constraints on the mixing angle, as given in table (2), the allowed parameter space for the perturbations $[\varepsilon_{1},\varepsilon_{2}]$ are obtained as shown in Fig.
-(0(a)).
+(1(a)).
 We use the same set of mixing angle constraints along with the bounds on two mass squared differences to assess the effective neutrino mass, which is given by
@@
 In Fig.
-(0(b)) we show the dependence of $|m_{ee}|$ on the phase $\beta$.
-The plot shows that the allowed effective mass in turn has imposed a bound on the phase $\beta\leq 270^{\circ}$.
-Apart from the bound on the effective neutrino mass $|m_{ee}|\lesssim 0.15$ eV, the bound on the total neutrino mass $\Sigma=m_{1}+m_{2}+m_{3}<0.12$ eV [6] puts a more stringent restriction on the parameter space of phase $\beta$ as $\sim[120^{\circ},240^{\circ}]$.
+(1(b)) we show the dependence of $|m_{ee}|$ on the phase $\beta$.
+In generating the plot using Eq.(2.36), we have calculated the masses, $m_{1},m_{2}$ and $m_{3}$ using Eq.(2.29), which requires the use of equations (2.27) and (2.28).
+Also, we ensure the values of $z_{a}$ to be positive as per Eq.(2.20).
+The plot shows that the allowed effective mass in turn has imposed a bound on the phase $\beta$ $\sim(100^{\circ}-260^{\circ})$.
 We have used this restricted parameter space of the phase $\beta$ to calculate the flavored CP asymmetries in section (4).
@@
 To estimate the CP asymmetry from the model, we have chosen the small perturbation parameters from the allowed parameter space shown in Fig.(1(a)), as $\varepsilon_{1}\sim-0.125$ and $\varepsilon_{2}\sim 0.075$.
-We have varied the phase $\beta$ within the range $[120^{\circ},240^{\circ}]$ so that it also remains consistent with the latest cosmological bound on the sum of the neutrino masses and the effective neutrino mass $|m_{ee}|\lesssim 0.15$ eV.
+We have varied the phase $\beta$ within the range $(100^{\circ}-260^{\circ})$ so that it remains consistent with the latest bound on the effective neutrino mass $|m_{ee}|\lesssim 0.15$ eV.
 The phase $\sigma_{1}$ is taken in such a way that it satisfies the relations given in Eq.(2.24).
-The viable values of the parameters $z_{1}$ and $z_{2}$ are calculated as functions of two mass squared differences and phase $\beta$, from the expressions equations (2.22) and (2.23) by making sure that it is consistent with the bound on total neutrino mass $\Sigma=m_{1}+m_{2}+m_{3}<0.12$ eV through the expression in Eq.(2.24).
+The viable values of the parameters $z_{1}$ and $z_{2}$ are calculated as functions of two mass squared differences and phase $\beta$, from the expressions equations (2.27) and (2.28).
 It is observed that CP asymmetries of the order $\mathcal{O}(10^{-7})$ can be produced for triplet masses $\sim 10^{10}$ GeV.
@@
 This feature also enhances the efficiency of leptogenesis which can be seen from the increase in the final baryon asymmetry as shown in the table.
-$M_{1}({\rm GeV})$
-$B^{L}_{1}$
-$\beta(^{\circ})$
-$\epsilon^{ee}_{1}/10^{-8}$
-$\epsilon^{e\mu}_{1}/10^{-9}$
-$\epsilon^{\mu\mu}_{1}/10^{-9}$
-$\epsilon^{\tau\tau}_{1}/10^{-9}$
-$|\eta_{B}|/10^{-10}$
-$0.48$
-$178$
-$0.70$
-$0.84$
-$1.01$
-$-1.68$
-$0.04$
-$3\times 10^{10}$
-$0.06$
-$175$
-$2.06$
-$2.46$
-$2.94$
-$-4.91$
-$0.29$
-$0.005$
-$165$
-$2.04$
-$2.43$
-$2.91$
-$-4.86$
-$0.73$
-$0.0005$
-$126$
-$2.06$
-$2.46$
-$2.94$
-$-4.91$
-$1.74$
-$0.34$
-$178$
-$2.79$
-$3.32$
-$3.98$
-$-6.63$
-$0.30$
-$4\times 10^{10}$
-$0.05$
-$175$
-$2.82$
-$3.36$
-$4.03$
-$-6.71$
-$0.44$
-$0.005$
-$164$
-$2.82$
-$3.36$
-$4.03$
-$-6.71$
-$1.07$
-$0.0006$
-$126$
-$2.83$
-$3.37$
-$4.04$
-$-6.74$
-$2.50$
+$M_{1}({\rm GeV})$ $B^{L}_{1}$ $\beta(^{\circ})$ $\epsilon^{ee}_{1}/10^{-8}$ $\epsilon^{e\mu}_{1}/10^{-9}$ $\epsilon^{\mu\mu}_{1}/10^{-9}$ $\epsilon^{\tau\tau}_{1}/10^{-9}$ $|\eta_{B}|/10^{-10}$
+$0.48$ $178$ $0.70$ $0.84$ $1.01$ $-1.68$ $0.04$
+$3\times 10^{10}$ $0.06$ $175$ $2.06$ $2.46$ $2.94$ $-4.91$ $0.29$
+$0.005$ $165$ $2.04$ $2.43$ $2.91$ $-4.86$ $0.73$
+$0.0005$ $126$ $2.06$ $2.46$ $2.94$ $-4.91$ $1.74$
+$0.34$ $178$ $2.79$ $3.32$ $3.98$ $-6.63$ $0.30$
+$4\times 10^{10}$ $0.05$ $175$ $2.82$ $3.36$ $4.03$ $-6.71$ $0.44$
+$0.005$ $164$ $2.82$ $3.36$ $4.03$ $-6.71$ $1.07$
+$0.0006$ $126$ $2.83$ $3.37$ $4.04$ $-6.74$ $2.50$
 Table 3: Baryon asymmetries through flavored leptogenesis for the triplet mass hierarchy case-1, $M_{1}\ll M_{2}$ by solving DME in the energy range $[10^{9},10^{12}]$ GeV for different values of $M_{1}\sim 10^{10}$ GeV, $\epsilon^{\alpha\beta}_{1}$, and hierarchical $(B^{L}_{1}\sim 0.0005,0.005,0.05)$ as well as comparable $(B^{L}_{1}\sim 0.5)$ branching ratio $B^{L}_{1}$.
@@
 We have plotted the obtained flavored lepton asymmetries and baryon asymmetry in Fig.
-(3(a)) along with the evolution of different washout terms in Fig.
-(3(b)).
+(4(a)) along with the evolution of different washout terms in Fig.
+(4(b)).
 Although we have shown the results for final baryon asymmetry for a few benchmark points, we have taken different data sets corresponding to particular flavored CP asymmetries for $M_{1}\sim 10^{10}$ GeV and hierarchical branching ratios and we have plotted the obtained baryon asymmetry for each set in Fig.
-(5(a)).
-$M_{1}({\rm GeV})$
-$B^{L}_{1}$
-$|u_{1}|(\rm eV)$
-$\epsilon^{ee}_{1}/10^{-7}$
-$\epsilon^{e\mu}_{1}/10^{-8}$
-$\epsilon^{\mu\mu}_{1}/10^{-8}$
-$\epsilon^{\tau\tau}_{1}/10^{-8}$
-$|\eta_{B}|/10^{-10}$
-$3\times 10^{10}$ Fig. (4)
-$0.58$
-$4.80$
-$4.47$
-$5.32$
-$6.38$
-$-1.06$
-$5.29$
+(6(a)).
+$M_{1}({\rm GeV})$ $B^{L}_{1}$ $|u_{1}|(\rm eV)$ $\epsilon^{ee}_{1}/10^{-7}$ $\epsilon^{e\mu}_{1}/10^{-8}$ $\epsilon^{\mu\mu}_{1}/10^{-8}$ $\epsilon^{\tau\tau}_{1}/10^{-8}$ $|\eta_{B}|/10^{-10}$
+$3\times 10^{10}$ Fig. (4) $0.58$ $4.80$ $4.47$ $5.32$ $6.38$ $-1.06$ $5.29$
 Table 4: Baryon asymmetries through leptogenesis by solving DME in the energy range $[10^{9},10^{12}]$ GeV for $M_{1}=3\times 10^{10}$ GeV.
@@
 We have plotted the obtained flavored lepton asymmetries and baryon asymmetry in Fig.
-(4(a)) along with the evolution of different washout terms in Fig.
-(4(b)).
+(5(a)) along with the evolution of different washout terms in Fig.
+(5(b)).
 Similar to the case -I, we have taken different data sets corresponding to particular flavored CP asymmetries for $M_{2}\sim 10^{10}$ GeV and hierarchical branching ratios and we have plotted the obtained baryon asymmetry for each set in Fig.
-(5(b)).
+(6(b)).
 Figure 5: Solution of DMEs for asymmetries and washouts for set with $M_{2}=1.4\times 10^{10}$ GeV by using the flavored CP asymmetries from table (6).
@@
 Right: The color lines represent numerically obtained amount of washouts along different flavor directions in units of $sHz$.
-$M_{2}({\rm GeV})$
-$B^{L}_{2}$
-$\beta(^{\circ})$
-$\epsilon^{ee}_{2}/10^{-8}$
-$\epsilon^{e\mu}_{2}/10^{-9}$
-$\epsilon^{\mu\mu}_{2}/10^{-9}$
-$\epsilon^{\tau\tau}_{2}/10^{-7}$
-$|\eta_{B}|/10^{-10}$
-$0.59$
-$176$
-$-0.66$
-$-0.79$
-$-0.95$
-$0.75$
-$0.04$
-$1\times 10^{10}$
-$0.06$
-$172$
-$-0.67$
-$-0.80$
-$-0.96$
-$0.76$
-$0.07$
-$0.005$
-$157$
-$-0.67$
-$-0.80$
-$-0.96$
-$0.75$
-$0.86$
-$0.0006$
-$126$
-$-0.73$
-$-0.87$
-$-1.05$
-$0.75$
-$3.26$
-$0.53$
-$177$
-$-1.29$
-$-1.53$
-$-1.84$
-$1.45$
-$0.01$
-$2\times 10^{10}$
-$0.05$
-$176$
-$-1.42$
-$-1.68$
-$-2.02$
-$1.59$
-$0.03$
-$0.006$
-$156$
-$-1.42$
-$-1.69$
-$-2.03$
-$1.58$
-$1.66$
-$0.0005$
-$126$
-$-1.42$
-$-1.69$
-$-2.03$
-$1.45$
-$7.51$
+$M_{2}({\rm GeV})$ $B^{L}_{2}$ $\beta(^{\circ})$ $\epsilon^{ee}_{2}/10^{-8}$ $\epsilon^{e\mu}_{2}/10^{-9}$ $\epsilon^{\mu\mu}_{2}/10^{-9}$ $\epsilon^{\tau\tau}_{2}/10^{-7}$ $|\eta_{B}|/10^{-10}$
+$0.59$ $176$ $-0.66$ $-0.79$ $-0.95$ $0.75$ $0.04$
+$1\times 10^{10}$ $0.06$ $172$ $-0.67$ $-0.80$ $-0.96$ $0.76$ $0.07$
+$0.005$ $157$ $-0.67$ $-0.80$ $-0.96$ $0.75$ $0.86$
+$0.0006$ $126$ $-0.73$ $-0.87$ $-1.05$ $0.75$ $3.26$
+$0.53$ $177$ $-1.29$ $-1.53$ $-1.84$ $1.45$ $0.01$
+$2\times 10^{10}$ $0.05$ $176$ $-1.42$ $-1.68$ $-2.02$ $1.59$ $0.03$
+$0.006$ $156$ $-1.42$ $-1.69$ $-2.03$ $1.58$ $1.66$
+$0.0005$ $126$ $-1.42$ $-1.69$ $-2.03$ $1.45$ $7.51$
 Table 5: Baryon asymmetries through flavored leptogenesis for the triplet mass hierarchy case - II, $M_{2}\ll M_{1}$ by solving DME in the energy range $[10^{9},10^{12}]$ GeV for different values of $M_{2}\sim 10^{10}$ GeV, $\epsilon^{\alpha\beta}_{2}$, and hierarchical $(B^{L}_{2}\sim 0.0005,0.005,0.05)$ as well as comparable $(B^{L}_{2}\sim 0.5)$ branching ratio $B^{L}_{2}$.
-$M_{2}({\rm GeV})$
-$B^{L}_{2}$
-$|u_{2}|(\rm eV)$
-$\epsilon^{ee}_{2}/10^{-8}$
-$\epsilon^{e\mu}_{2}/10^{-9}$
-$\epsilon^{\mu\mu}_{2}/10^{-9}$
-$\epsilon^{\tau\tau}_{2}/10^{-7}$
-$|\eta_{B}|/10^{-10}$
-$1.4\times 10^{10}$ Fig. (5)
-$0.0006$
-$48.1$
-$-1.00$
-$-1.20$
-$-1.44$
-$1.04$
-$4.92$
+$M_{2}({\rm GeV})$ $B^{L}_{2}$ $|u_{2}|(\rm eV)$ $\epsilon^{ee}_{2}/10^{-8}$ $\epsilon^{e\mu}_{2}/10^{-9}$ $\epsilon^{\mu\mu}_{2}/10^{-9}$ $\epsilon^{\tau\tau}_{2}/10^{-7}$ $|\eta_{B}|/10^{-10}$
+$1.4\times 10^{10}$ Fig. (5) $0.0006$ $48.1$ $-1.00$ $-1.20$ $-1.44$ $1.04$ $4.92$
 Table 6: Baryon asymmetries through leptogenesis by solving DME in the energy range $[10^{9},10^{12}]$ GeV for $M_{2}=1.4\times 10^{10}$ GeV.
@@
 The amount of washouts greatly influences the efficiency of leptogenesis.
-Comparing figures (3(b)) ($B_{1}^{L}=0.58$) and (4(b)) ($B_{2}^{L}=0.0006$), one can see that in the first case, the washout effects along the flavor directions are more as compared to the second case (except $ee$-flavor direction).
+Comparing figures (4(b)) ($B_{1}^{L}=0.58$) and (5(b)) ($B_{2}^{L}=0.0006$), one can see that in the first case, the washout effects along the flavor directions are more as compared to the second case (except $ee$-flavor direction).
 This feature can be attributed to the specific flavor structure of the triplet Yukawa coupling matrices of the model.
@@
 (4), also Fig.
-(5(a))) the observable range of baryon asymmetry can be achieved for triplet mass $M_{1}$ as low as $\sim 3\times 10^{10}$ GeV for comparable branching ratios $B^{L}_{1}\sim 0.5$.
+(6(a))) the observable range of baryon asymmetry can be achieved for triplet mass $M_{1}$ as low as $\sim 3\times 10^{10}$ GeV for comparable branching ratios $B^{L}_{1}\sim 0.5$.
 Taking hierarchical branching ratios and requiring adequate flavored CP asymmetries, baryon asymmetry can only be produced in the observable range for a triplet mass $M_{1}>5\times 10^{10}$ GeV for a highly hierarchical branching ratio $B^{L}_{1}\sim 0.0005$.
@@
 5 ( also Fig.
-(5(b))) the observational range of baryon asymmetry can be achieved for triplet mass $M_{2}$ as low as $\sim 1\times 10^{10}$ GeV for highly hierarchical branching ratio $B^{L}_{2}\sim 0.0005$.
+(6(b))) the observational range of baryon asymmetry can be achieved for triplet mass $M_{2}$ as low as $\sim 1\times 10^{10}$ GeV for highly hierarchical branching ratio $B^{L}_{2}\sim 0.0005$.
     3. (c)
@@
 As shown in tables (3) and (5), the suppression in CP asymmetry due to the hierarchical branching ratio, can be compensated by the extra source of CP violation due to the presence of the SCPV phase $\beta$ in the neutrino mass matrix.
-As a result, even with the modified washout effect from hierarchical branching ratios, it is possible to achieve maximal CP violation and baryon asymmetry can be achieved in the observable range.
+As a result, even with the modified washout effect from hierarchical branching ratios, it is possible to achieve maximal CP violation and baryon asymmetry in the observable range.
 The flavor structure of the model is suitable to draw these conclusions.
@@
 In this paper, we have studied, in an extended Standard model, neutrino mass generated via type-II seesaw mechanism and leptogenesis from the decay of heavy scalar triplet with a common source of CP violation.
-The model provides a viable explanation of baryon asymmetry through thermal leptogenesis, by sourcing enough CP asymmetry spontaneously broken due to the presence of a scalar singlet with complex VEV.
+The model provides a viable explanation of baryon asymmetry through thermal leptogenesis, by sourcing enough CP symmetry spontaneously broken due to the presence of a scalar singlet with complex VEV.
 Moreover, the neutrino mass matrix exhibits TBM structure which generally fails to address the occurrence of non-zero CHOOZ mixing angle, $\theta_{13}$.
 By incorporating small perturbations around flavon VEV in the high-energy scale, it can accommodate $\theta_{13}\neq 0$ in the model.
-With the latest neutrino oscillation data, bound on total neutrino mass and the effective neutrino mass, we have analyzed the parameter space of the model.
+With the latest neutrino oscillation data and bound on the effective neutrino mass, we have analyzed the parameter space of the model.
 Using the constrained parameter space of the model we have studied flavored leptogenesis from the decay of the triplet scalars.
@@
 Finally from equations (4.1), (C.1), (C.2), (C.3) and (C.4), the flavored CP asymmetries $\epsilon^{\alpha\beta}_{1}$ are formed as,
-$\epsilon^{ee}_{1}=\frac{M^{2}_{1}|u_{1}|^{2}}{2\pi\left(3z^{2}_{1}v^{4}+4M^{4}_{1}|u_{1}|^{4}\right)}\Big{[}\frac{2}{3}z_{1}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\left(1-\varepsilon_{1}-\varepsilon_{2}\right)\sin(\beta-\sigma_{1})\Big{]},$ (C.8)
-$\epsilon^{e\mu}_{1}$$\approx$ $\frac{M^{2}_{1}|u_{1}|^{2}}{2\pi\left(3z^{2}_{1}v^{4}+4M^{4}_{1}|u_{1}|^{4}\right)}\Big{[}-\frac{2}{3}\varepsilon_{1}z_{1}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})\Big{]}$ (C.9)
+$\epsilon^{ee}_{1}=\frac{M_{1}(B^{L}_{1}B^{\phi}_{1})^{\frac{1}{2}}c_{ee}}{v^{2}[{\rm Tr}(m^{(1)\dagger}_{\nu}m^{(1)}_{\nu})]^{1/2}}\Big{[}\frac{2}{3}z_{1}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\left(1-\varepsilon_{1}-\varepsilon_{2}\right)\sin(\beta-\sigma_{1})\Big{]},$ (C.8)
+$\epsilon^{e\mu}_{1}$$\approx$ $\frac{M_{1}(B^{L}_{1}B^{\phi}_{1})^{\frac{1}{2}}c_{e\mu}}{v^{2}[{\rm Tr}(m^{(1)\dagger}_{\nu}m^{(1)}_{\nu})]^{1/2}}\Big{[}-\frac{2}{3}\varepsilon_{1}z_{1}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})\Big{]}$ (C.9)
 $\approx$ $\epsilon^{\mu e}_{1},$
-$\epsilon^{\mu\mu}_{1}$$\approx$ $\frac{M^{2}_{1}|u_{1}|^{2}}{2\pi\left(3z^{2}_{1}v^{4}+4M^{4}_{1}|u_{1}|^{4}\right)}\times\varepsilon_{2}z_{1}\Big{[}\frac{1}{3}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})$ (C.10)
+$\epsilon^{\mu\mu}_{1}$$\approx$ $\frac{M_{1}(B^{L}_{1}B^{\phi}_{1})^{\frac{1}{2}}c_{\mu\mu}}{v^{2}[{\rm Tr}(m^{(1)\dagger}_{\nu}m^{(1)}_{\nu})]^{1/2}}\times\varepsilon_{2}z_{1}\Big{[}\frac{1}{3}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})$ (C.10)
 $+$ $\left(z^{2}_{1}+z^{2}_{2}-2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{2})\Big{]},$
 and
-$\epsilon^{\tau\tau}_{1}$$\approx$ $\frac{M^{2}_{1}|u_{1}|^{2}}{2\pi\left(3z^{2}_{1}v^{4}+4M^{4}_{1}|u_{1}|^{4}\right)}\times\varepsilon_{1}z_{1}\Big{[}\frac{1}{3}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})$ (C.11)
+$\epsilon^{\tau\tau}_{1}$$\approx$ $\frac{M_{1}(B^{L}_{1}B^{\phi}_{1})^{\frac{1}{2}}c_{\tau\tau}}{v^{2}[{\rm Tr}(m^{(1)\dagger}_{\nu}m^{(1)}_{\nu})]^{1/2}}\times\varepsilon_{1}z_{1}\Big{[}\frac{1}{3}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})$ (C.11)
 $+$ $\left(z^{2}_{1}+z^{2}_{2}-2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{2})\Big{]}.$
 From equations (4.2), (C.5), (C), (C.6) and (C.7), the flavored CP asymmetries $\epsilon^{\alpha\beta}_{2}$ are formed as,
-$\epsilon^{ee}_{2}$ $=$ $\frac{M^{2}_{2}|u_{2}|^{2}}{4\pi\left(z^{2}_{2}v^{4}(1-2\varepsilon_{1}-2\varepsilon_{2})+2M^{4}_{2}|u_{2}|^{4}\right)}\times-\frac{2}{3}z_{2}$ (C.12)
+$\epsilon^{ee}_{2}$ $=$ $\frac{M_{2}(B^{L}_{2}B^{\phi}_{2})^{\frac{1}{2}}c_{ee}}{v^{2}[{\rm Tr}(m^{(2)\dagger}_{\nu}m^{(2)}_{\nu})]^{1/2}}\times-\frac{2}{3}z_{2}\times$ (C.12)
 $\Big{[}\frac{2}{3}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}(1-2\varepsilon_{1}-2\varepsilon_{2})\sin\sigma_{1}$
 $+$ $\frac{1}{3}z_{1}(1+\varepsilon_{1}+\varepsilon_{2})\sin\beta\Big{]},$
-$\epsilon^{e\mu}_{2}$$\approx$ $\frac{M^{2}_{2}|u_{2}|^{2}}{4\pi\left(z^{2}_{2}v^{4}(1-2\varepsilon_{1}-2\varepsilon_{2})+2M^{4}_{2}|u_{2}|^{4}\right)}\times\frac{1}{3}z_{2}$ (C.13)
+$\epsilon^{e\mu}_{2}$$\approx$ $\frac{M_{2}(B^{L}_{2}B^{\phi}_{2})^{\frac{1}{2}}c_{e\mu}}{v^{2}[{\rm Tr}(m^{(2)\dagger}_{\nu}m^{(2)}_{\nu})]^{1/2}}\times\frac{1}{3}z_{2}\times$ (C.13)
 $\Big{[}(\frac{7}{6}\varepsilon_{1}+\frac{1}{6}\varepsilon_{2}-\frac{1}{3})\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin\sigma_{1}$
@@
 $\approx$ $\epsilon^{\mu e}_{2},$
-$\epsilon^{\mu\mu}_{2}$$\approx$ $\frac{M^{2}_{2}|u_{2}|^{2}}{4\pi\left(z^{2}_{2}v^{4}(1-2\varepsilon_{1}-2\varepsilon_{2})+2M^{4}_{2}|u_{2}|^{4}\right)}\times-\frac{2}{3}z_{2}$ (C.14)
+$\epsilon^{\mu\mu}_{2}$$\approx$ $\frac{M_{2}(B^{L}_{2}B^{\phi}_{2})^{\frac{1}{2}}c_{\mu\mu}}{v^{2}[{\rm Tr}(m^{(2)\dagger}_{\nu}m^{(2)}_{\nu})]^{1/2}}\times-\frac{2}{3}z_{2}\times$ (C.14)
 $\Big{[}(-\frac{5}{6}\varepsilon_{1}+\frac{1}{6}\varepsilon_{2}-\frac{1}{6})\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin\sigma_{1}$
@@
 and
-$\epsilon^{\tau\tau}_{2}$$\approx$ $\frac{M^{2}_{2}|u_{2}|^{2}}{4\pi\left(z^{2}_{2}v^{4}(1-2\varepsilon_{1}-2\varepsilon_{2})+2M^{4}_{2}|u_{2}|^{4}\right)}\times-\frac{2}{3}z_{2}$ (C.15)
+$\epsilon^{\tau\tau}_{2}$$\approx$ $\frac{M_{2}(B^{L}_{2}B^{\phi}_{2})^{\frac{1}{2}}c_{\tau\tau}}{v^{2}[{\rm Tr}(m^{(2)\dagger}_{\nu}m^{(2)}_{\nu})]^{1/2}}\times-\frac{2}{3}z_{2}\times$ (C.15)
 $\Big{[}(\frac{1}{6}\varepsilon_{1}-\frac{5}{6}\varepsilon_{2}+\frac{1}{6})\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin\sigma_{1}$